### v4

# latexOverwriting file

Abstract
The ensemble Kalman filter is widely used in applications because, for high dimensional filtering problems, it has a robustness that is not shared for example by the particle filter; in particular it does not suffer from weight collapse. However, there is no theory which quantifies its accuracy as an approximation of the true filtering distribution, except in the Gaussian setting. To address this issue we provide the first analysis of the accuracy of the ensemble Kalman filter beyond the Gaussian setting. We prove two types of results: the first type comprise a stability estimate controlling the error made by the ensemble Kalman filter in terms of the difference between the true filtering distribution and a nearby Gaussian; and the second type use this stability result to show that, in a neighbourhood of Gaussian problems, the ensemble Kalman filter makes a small error, in comparison with the true filtering distribution. Our analysis is developed for the mean field ensemble Kalman filter. We rewrite the update equations for this filter, and for the true filtering distribution, in terms of maps on probability measures. We introduce a weighted total variation metric to estimate the distance between the two filters and we prove various stability estimates for the maps defining the evolution of the two filters, in this metric. Using these stability estimates we prove results of the first and second type, in the weighted total variation metric. We also provide a generalization of these results to the Gaussian projected filter, which can be viewed as a mean field description of the unscented Kalman filter.
Keywords. Ensemble Kalman filter, Stochastic filtering, Weighted total variation metric, Stability estimates, Accuracy estimates, Near-Gaussian setting.
AMS subject classification. 60G35, 62F15, 65C35, 70F45, 93E11.
1 Introduction
1.1 Context
This paper is concerned with the study of partially and noisily observed dynamical systems. Filtering refers to the sequential updating of the probability distribution of the state of the (possibly stochastic) dynamical system, given partial noisy observations [3, 43, 51]. The Kalman filter, introduced in 1960, provides an explicit solution to this problem in the setting of linear dynamical systems, linear observations and additive Gaussian noise [38]; the desired probability distribution is Gaussian and the Kalman filter provides explicit update formulae for the mean and covariance. The extended Kalman filter is a linearization-based methodology which was developed in the 1960s and 1970s to apply to situations beyond the linear-Gaussian setting [36]. It is, however, not practical in high dimensions, because of the need to compute, and sequentially update, large covariance matrices [29]. To address this issue, the ensemble Kalman filter was introduced in 1994 [25], using ensemble-based low-rank approximations of the covariances, and is a method well-adapted to high dimensional problems. The unscented Kalman filter, introduced in 1997, provides an alternative approach to nonlinear and non-Gaussian problems [37], using quadrature to approximate covariance matrices, and is well-adapted to problems of moderate dimension. The particle filter [20] is a provably convergent methodology for approximating the filtering distribution [11, 50]. However it does not scale well to high dimensional problems [4, 56]; this motivates the increasing adoption of ensemble Kalman methods.
Over the last two decades ensemble Kalman filters have found widespread use in the geophysical sciences, are starting to be used in other application domains, and have been developed as a general purpose tool for solving inverse problems; for reviews of such developments see [8, 24, 26]. Despite widespread and growing adoption, theory quantifying the accuracy of the ensemble Kalman filter, in relation to the true filtering distribution, is limited. Two exceptions are the important contributions [44, 46] which concern accuracy in the large particle limit in the setting where the underlying filtering problem is Gaussian. However, there is no proof that the ensemble Kalman methodology can accurately approximate the desired filtering distribution beyond the Gaussian setting. Indeed, in general the methodology does not correctly reproduce the filtering distribution; for examples and analysis in the setting of filtering and Bayesian inverse problems see [1, 44] and [22] respectively. The aim of our work is to address this issue by proving accuracy of the ensemble Kalman filter beyond the Gaussian setting; specifically our analysis applies when the true filtering distribution is close to Gaussian, after appropriate lifting to the joint space of state and data. We perform the analysis for the mean field ensemble Kalman filter, focusing on quantifying the effect of the Gaussian approximation underlying the ensemble Kalman filter. We also study the Gaussian projected filter; this filter can be seen as a mean field version of the unscented Kalman filter. Both the ensemble and unscented Kalman mean field models are defined in [8].
The three primary contributions of our paper are: (i) to establish finite time stability estimates, in an appropriate weighted TV metric which enables control of first and second order moments, for various nonlinear maps required to define the nonlinear Markov processes determining filter evolution; (ii) to use these results to establish stability estimates controlling the error made by the mean field ensemble Kalman filter in terms of the difference between the true filtering distribution and a nearby Gaussian; and (iii) to deploy these estimates to prove closeness of the mean field ensemble Kalman filters and the true filtering distribution, in the near-Gaussian setting. These results are also established for the mean-field unscented Kalman filter. As well as making the three primary contributions, the work suggests many questions for further analysis and the numerical analysis framework we deploy (“consistency plus stability implies convergence”) is a natural one in which to pursue these questions. In particular we assume bounded vector fields and discuss only finite time error; overcoming these assumptions requires new ideas and is outside the scope of this first paper.
Although our study of the accuracy of ensemble Kalman filters, beyond the Gaussian setting, is new, there exists a growing body of literature analyzing ensemble Kalman methods from different perspectives. In the context of long-time behaviour see [39, 60, 13, 12]. For filter divergence see [30, 40]. For localization see [59, 58]. For high dimensions see [45]. For continuous-time and mean-field limits see [41, 42, 23]. In the context of inverse problems see [35, 54, 55, 34, 33].
1.2 Overview of Paper
In Subsection 1.3 which follows we define useful notation employed throughout the paper. Then, in Subsection 1.4 we define the filtering problem as employed throughout the paper, building on this notation. Section 2 introduces the mean field ensemble Kalman filter and proves our main approximation theorem, in the near Gaussian setting; the theorem is based on a stability estimate which transfers distance between the true filter and its Gaussian projection into the distance between the true filter and the ensemble Kalman filter. In Section 3 we define, and then state and prove related theory for, the Gaussian projected filter. Although we provide theorems of a type not seen before for nonlinear Kalman filters, and new methods of analysis to derive them, our work leads to many substantial open problems; we conclude in Section 4 by highlighting those we have identified as being of particular value to advancing the field. The reader may wish to study the concluding Section 4, in conjunction with our set-up in Section 2, to understand the specific problem formulation for our main theorems and to appreciate the substantial challenges to be met in order to build on, and go beyond, our theorems. The key auxiliary results underpinning the proof of our main theorems are given in Appendices B and C; these rely on technical results presented in Appendix A.
1.3 Notation
[…]
$\mathsf{G}\mu=\mathcal{N}\bigl{(}\mathcal{M}(\mu),\mathcal{C}(\mu)\bigr{)}.$
We observe [5] that
$\mathsf{G}\mu=\operatorname*{arg\,min}_{\nu\in\mathcal{G}}\mathrm{KL}(\mu\|\nu),$
where $\mathrm{KL}(\mu\|\nu)$ is the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$, defined in (B.46). Note that $\mathsf{G}$ defines a nonlinear mapping. We refer to $\mathsf{G}$ as a projection because of its characterization as finding closest point to $\mu$ with respect to $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}).$ Throughout this paper all probability measures have Lebesgue density, because of our assumptions concerning the noise structure in the dynamics model and the data acquisition model. Thus we abuse notation by using the same symbols for probability measures and their densities. For a probability measure $\mu\in\mathcal{P}(^{n})$, the notation $\mu(u)$ for $u\in^{n}$ refers to the Lebesgue density of $\mu$ evaluated at $u$, whereas $\mu[f]$ for a function $f\colon^{n}\to$ is a short-hand notation for $\int_{{}^{n}}f\mathrm{d}\mu$.
The notations $\mathcal{M}(\mu)$ and $\mathcal{C}(\mu)$ denote respectively the mean and covariance under $\mu$:
$\mathcal{M}(\mu)=\mu[u],\qquad\mathcal{C}(\mu)=\mu\Bigl{[}\bigl{(}u-\mathcal{M}(\mu)\bigr{)}\otimes\bigl{(}u-\mathcal{M}(\mu)\bigr{)}\Bigr{]}.$
The notation $\mathcal{P}_{R}(^{n})$ for $R\geqslant 1$ refers to the subset of $\mathcal{P}(^{n})$ of probability measures whose first and second moments satisfy the bound
$\lvert\mathcal{M}(\mu)\rvert\leqslant R,\qquad\frac{1}{R^{2}}\mathit{I}_{n}\preccurlyeq\mathcal{C}(\mu)\preccurlyeq R^{2}\mathit{I}_{n}.$ (1.1)
[…]
For a probability measure $\pi\in\mathcal{P}(^{d}\times^{K})$ associated with random variable $(u,y)\in^{d}\times^{K}$ we use the notation $\mathcal{M}^{u}(\pi)\in^{d}$, $\mathcal{M}^{y}(\pi)\in^{K}$ for the means of the marginal distributions, and the notation $\mathcal{C}^{uu}(\pi)\in^{d\times d}$, $\mathcal{C}^{uy}(\pi)\in^{d\times K}$ and $\mathcal{C}^{yy}(\pi)\in^{K\times K}$ for the blocks of the covariance matrix $\mathcal{C}(\pi)$. That is to say,
$\mathcal{M}(\pi)=\begin{pmatrix}\mathcal{M}^{u}(\pi)\\ \mathcal{M}^{y}(\pi)\end{pmatrix},\qquad\mathcal{C}(\pi)=\begin{pmatrix}\mathcal{C}^{uu}(\pi)&\mathcal{C}^{uy}(\pi)\\ \mathcal{C}^{uy}(\pi)^{\mathsf{T}}&\mathcal{C}^{yy}(\pi)\end{pmatrix}.$
Throughout this work, we employ the following weighted total variation distance.
Definition 1.
Let $g\colon^{n}\to[1,\infty)$ be given by $g(v)=1+\lvert v\rvert^{2}$. We define the weighted total variation metric $d_{g}\colon\mathcal{P}(^{n})\times\mathcal{P}(^{n})\to$ by
$d_{g}(\mu_{1},\mu_{2})=\sup_{\lvert f\rvert\leqslant g}\bigl{\lvert}\mu_{1}[f]-\mu_{2}[f]\bigr{\rvert},$
where the supremum is over all functions $f\colon^{n}\to$ which are bounded from above by $g$ pointwise and in absolute value.
Metrics of this type have been employed previously in the literature. See for example the early reference [48], where a weighted total variation metric is used for stuyding ergodicity for Markov chains, and the reference [47], where a similar metric is employed in the context of perturbation of Markov chains. See also [32], where a direct proof of Harris’ ergodic theorem relying on appropriate weighted total variation norms is given, as well as [63, Theorem 6.15], which states that Wasserstein distances can be controlled by weighted total variation metrics. There are other dual metrics on probability measures commonly employed in the literature, such as the so-called maximum mean discrepancy, which is based on a dual formulation in a reproducing kernel Hilbert space; see e.g. [57].
Remark 2.
If $\mu_{1},\mu_{2}$ have Lebesgue densities $\rho_{1}$, $\rho_{2}$, then
$d_{g}(\mu_{1},\mu_{2})=\int g(v)\left\lvert\rho_{1}(v)-\rho_{2}(v)\right\rvert\,\mathrm{d}v.$
Remark 3.
Unlike the usual total variation distance, the weighted metric in Definition 1 enables control of the differences $\lvert\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\rvert$ and $\lVert\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\rVert$; several lemmas used to prove our main results rely on this control. More precisely, it is possible to show that for all $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ with finite second moments, it holds that
$\bigl{\lvert}\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\bigr{\rvert}$ $\leqslant\frac{1}{2}d_{g}(\mu_{1},\mu_{2}),$
$\bigl{\lVert}\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\bigr{\rVert}$ $\leqslant\left(1+\frac{1}{2}\lvert\mathcal{M}(\mu_{1})+\mathcal{M}(\mu_{2})\rvert\right)\,d_{g}(\mu_{1},\mu_{2}).$
This is the content of Lemma 18, proved in the appendix.
1.4 Filtering Distribution
We consider a general setting in Subsection 1.4.1, and provide details for the Gaussian setting in particular in Subsection 1.4.2.
1.4.1 General Case
We consider the following stochastic dynamics and data model
$u_{j+1}$ $=\Psi(u_{j})+\xi_{j},\qquad$ $\xi_{j}\sim\mathcal{N}(0,\mathit{\Sigma}),$ (1.2a)
$y_{j+1}$ $=h(u_{j+1})+\eta_{j+1},\qquad$ $\eta_{j+1}\sim\mathcal{N}(0,\mathit{\Gamma}).$ (1.2b)
Here $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}$ is the unknown state, evolving in ${}^{d}$, and $\{y_{j}\}_{j\in\llbracket 0,J\rrbracket}$ are the observations, evolving in ${}^{K}$. We assume that the initial state is distributed according to a Gaussian distribution $u_{0}\sim\mathcal{N}(m_{0},\mathit{C}_{0})$ and that the following independence assumption is satisfied:
$u_{0}\perp\!\!\!\!\perp\xi_{0}\perp\!\!\!\!\perp\dotsb\perp\!\!\!\!\perp\xi_{J}\perp\!\!\!\!\perp\eta_{1}\perp\!\!\!\!\perp\dotsb\perp\!\!\!\!\perp\eta_{J+1}.$ (1.3)
The objective of probabilistic filtering is to sequentially estimate the conditional distribution of the unknown state, given the data, as new data arrives. The true filtering distribution $\mu_{j}$ is the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ of the data process up to step $j$. Data $Y_{j}$ may be thought of as arising from a realization of (1.4.1); but the case of model mis-specification, where $Y_{j}$ does not necessarily arise from (1.4.1), is also of interest.
It is well-known [43, 51] that the true filtering distribution evolves according to
$\mu_{j+1}=\mathsf{L}_{j}\mathsf{P}\mu_{j}.$ (1.4)
where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps on probability measures, referred to respectively as the prediction and analysis steps in the data assimilation community [3]. The operator $\mathsf{P}$ is linear and defined by the Markov kernel associated with the stochastic dynamics (1.2a). Its action on a probability measure with Lebesgue density $\mu$ reads
$\mathsf{P}\mu(u)=\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\mu(v)\,\mathrm{d}v.$ (1.5)
The operator $\mathsf{L}_{j}$ is a nonlinear map which formalizes the incorporation of the new datum $y^{\dagger}_{j+1}$ using Bayes’ theorem. Its action reads
$\mathsf{L}_{j}\mu(u)=\frac{\displaystyle\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)\mu(u)}{\displaystyle\int_{{}^{d}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)\mu(U)\,\mathrm{d}U}$ (1.6)
The operator $\mathsf{L}_{j}$ effects a reweighting of the measure to which it applies, with more weight assigned to the state values that are consistent with the observation. It is convenient in this work to decompose the analysis map $\mathsf{L}_{j}$ into the composition $\mathsf{B}_{j}\mathsf{Q}$, where the operators $\mathsf{Q}\colon\mathcal{P}(^{d})\to\mathcal{P}(^{d}\times^{K})$ and $\mathsf{B}_{j}\colon\mathcal{P}(^{d}\times^{K})\to\mathcal{P}(^{d})$ are defined by
$\mathsf{Q}\mu(u,y)=\frac{1}{\sqrt{(2\pi)^{K}\det\mathit{\Gamma}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)\mu(u).$ (1.7)
and
$\mathsf{B}_{j}\mu(u)=\frac{\mu(u,y_{j+1}^{\dagger})}{\displaystyle\int_{{}^{d}}\mu(U,y_{j+1}^{\dagger})\,\mathrm{d}U}.$ (1.8)
The operator $\mathsf{Q}$ maps a probability measure with density $\mu$ into the density associated with the joint random variable $\bigl{(}U,h(U)+\eta\bigr{)}$, where $U\sim\mu$ is independent of $\eta\sim\mathcal{N}(0,\mathit{\Gamma})$. The operator $\mathsf{B}_{j}$ performs conditioning on the data $y^{\dagger}_{j+1}$. Map $\mathsf{Q}$ is linear whilst $\mathsf{B}_{j}$ is nonlinear. We may thus write (1.4) in the form
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}.$ (1.9)
1.4.2 Gaussian Case
Now consider the setting in which $\Psi(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=M\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}$, and $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=H\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}$ for matrices $\mathit{M}\in^{d\times d}$ and $\mathit{H}\in^{K\times d}.$ In this case the filtering distribution is Gaussian: $\mu_{j}=\mathcal{N}(m_{j},C_{j}).$ The Kalman filter [38] gives explicit evolution equations for the pair $(m_{j},C_{j})\in^{d}\times^{d\times d}.$ To write these down it is helpful to note that $\widehat{\mu}_{j+1}:={\mathsf{P}}\mu_{j}$ is also Gaussian: $\widehat{\mu}_{j+1}=\mathcal{N}(\widehat{m}_{j},\widehat{\mathit{C}}_{j}).$
Then $(\mu_{j+1},\widehat{\mu}_{j+1})$ are determined from $\mu_{j}$ by the update formulae
$\widehat{m}_{j+1}$ $=\mathit{M}m_{j},$ (1.10a)
[…]
$m_{j+1}$ $=\widehat{m}_{j+1}+\widehat{\mathit{C}}_{j+1}\mathit{H}^{\mathsf{T}}\left(\mathit{H}\widehat{\mathit{C}}_{j+1}\mathit{H}^{\mathsf{T}}+\mathit{\Gamma}\right)^{-1}\left({y}^{\dagger}_{j+1}-\mathit{H}\widehat{m}_{j+1}\right),$ (1.10c)
$\mathit{C}_{j+1}$ $=\widehat{\mathit{C}}_{j+1}-\widehat{\mathit{C}}_{j+1}\mathit{H}^{\mathsf{T}}\left(\mathit{H}\widehat{\mathit{C}}_{j+1}\mathit{H}^{\mathsf{T}}+\mathit{\Gamma}\right)^{-1}\mathit{H}\widehat{\mathit{C}}_{j+1},$ (1.10d)
where $\{{y}^{\dagger}_{j}\}$ is the observation. Recall that this observation may be thought of as arising from a realization of (1.4.1); but, in the case of model mis-specification, may be derived from a different source.
2 The Ensemble Kalman Filter
Following from expository discussion of the Gaussian case in  Subsection 2.1, in Subsection 2.2 we define the specific version of the mean-field ensemble Kalman filter that we analyze here; other versions may be found in [8] and will be amenable to similar analyses. In Subsection 2.3 we prove our main stability theorem, showing that the error between the true filter and its Gaussian projection may be used to control the error between the true filter and the ensemble Kalman filter. In Subsection 2.4 we prove a corollary to this theorem, establishing that the mean field ensemble Kalman filter accurately approximates the true filter for a specific class of non-Gaussian problems.
2.1 The Algorithm: Gaussian Case
To motivate the mean-field ensemble Kalman filter we first consider the Gaussian case and introduce the stochastic dynamical system
$\widehat{v}_{j+1}$ $=Mv_{j}+\xi_{j},\qquad$ $\xi_{j}\sim\mathcal{N}(0,\mathit{\Sigma}),$ (2.1a)
$\widehat{y}_{j+1}$ $=H\widehat{v}_{j+1}+\eta_{j+1},\qquad$ $\eta_{j+1}\sim\mathcal{N}(0,\mathit{\Gamma})$ (2.1b)
$v_{j+1}$ $=\widehat{v}_{j+1}+\widehat{\mathit{C}}_{j+1}H^{\top}(H\widehat{\mathit{C}}_{j+1}H^{\top}+\mathit{\Gamma})^{-1}({y}^{\dagger}_{j+1}-\widehat{y}_{j+1}).$ (2.1c)
Here $\widehat{\mathit{C}}_{j+1}$ denotes the covariance of $\widehat{v}_{j+1}$. A simple calculation reveals that, if $v_{0}\sim\mathcal{N}(m_{0},\mathit{C}_{0})$, then in fact $\widehat{v}_{j+1}\sim\mathcal{N}(\widehat{m}_{j+1},\widehat{\mathit{C}}_{j+1})$ and $v_{j+1}\sim\mathcal{N}(m_{j+1},\mathit{C}_{j+1})$, where the means and covariances are given by the Kalman filter (1.4.2). Note that (2.1) constitutes a mean-field stochastic dynamical system because equation (2.1c) requires knowledge of $\widehat{\mathit{C}}_{j+1}$, which depends on the law $\mu_{j}$ of $v_{j}$. The law of this mean-field stochastic dynamical system is thus equal to the law of the Kalman filter. The analysis in [44, 46] is concerned with studying particle approximations of this Gaussian mean-field stochastic dynamical system, and convergence to the mean-field limit in the limit of infinite particles. In contrast, our work concerns the study of mean-field stochastic dynamical systems, but goes beyond the Gaussian setting. To this end, the next subsection defines the mean-field ensemble Kalman filter in the general, non-Gaussian, setting.
2.2 The Algorithm: General Case
The ensemble Kalman filter may be derived as a particle approximation of various mean field limits [8]. The specific mean field ensemble Kalman filter that we study in this paper reads
$\widehat{u}_{j+1}$ $=\Psi(u_{j})+\xi_{j},\qquad$ $\xi_{j}\sim\mathcal{N}(0,\mathit{\Sigma}),$ (2.2a)
$\widehat{y}_{j+1}$ $=h(\widehat{u}_{j+1})+\eta_{j+1},\qquad$ $\eta_{j+1}\sim\mathcal{N}(0,\mathit{\Gamma})$ (2.2b)
$u_{j+1}$ $=\widehat{u}_{j+1}+\mathcal{C}^{uy}\left(\widehat{\pi}_{j+1}^{K}\right)\mathcal{C}^{yy}\left(\widehat{\pi}_{j+1}^{K}\right)^{-1}\bigl{(}y^{\dagger}_{j+1}-\widehat{y}_{j+1}\bigr{)},$ (2.2c)
where $\widehat{\pi}^{K}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and independence of the noise terms (1.3) is still assumed to hold. See Subsection 1.3 for the definition of the covariance matrices that appear in (2.2c). Let us denote by $\mu^{K}_{j}$ the law of $u_{j}$. In order to rewrite the evolution of $\mu_{j}^{K}$ in terms of maps on probability measures, let us introduce for a given $y^{\dagger}_{j+1}$ the map $\mathsf{T}_{j}\colon\mathcal{P}(^{d}\times^{K})\to\mathcal{P}(^{d})$ defined by
$\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi$ (2.3)
Here the subscript ${}_{\sharp}$ denotes the pushforward and, for any given $\pi\in\mathcal{P}(^{d}\times^{K})$ and $z\in^{K}$, the map $\mathscr{T}$ is affine in its first two arguments and given by
$\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,z)\colon$${}^{d}\times^{K}\to^{d};$
$(u,y)\mapsto u+\mathcal{C}^{uy}(\pi)\mathcal{C}^{yy}(\pi)^{-1}\bigl{(}z-y\bigr{)}.$ (2.4)
As demonstrated in [8], with this notation the evolution of the probability measure $\mu_{j}^{K}$ may be written in compact form as
$\mu_{j+1}^{K}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{K}.$ (2.5)
We now discuss the preceding map in relation to (1.9). The specific affine map $\mathscr{T}$ used in (2.2c) is determined by the measure $\widehat{\pi}^{K}_{j+1}$ (here equal to $\mathsf{Q}\mathsf{P}\mu_{j}^{K}$) and the data $y^{\dagger}_{j+1}$. That the law of $u_{j}$ in (2.2) evolves according to (2.5) follows from the following observations:
  • If $u_{j}\sim\mu_{j}^{K}$, then $\widehat{u}_{j+1}\sim\mathsf{P}\mu_{j}^{K}$, by definition of $\mathsf{P}$.
  • Given the definition (1.7) of the operator $\mathsf{Q}$, the random vector $(\widehat{u}_{j+1},\widehat{y}_{j+1})$ is distributed according to $\widehat{\pi}^{K}_{j+1}=\mathsf{Q}\mathsf{P}\mu_{j}^{K}$.
  • Equation (2.2c) then implies that $u_{j+1}\sim\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{K}$.
As we show in Lemma 19, the operator $\mathsf{T}_{j}$ coincides with the conditioning operator $\mathsf{B}_{j}$ over the subset $\mathcal{G}(^{d}\times^{K})\subset\mathcal{P}(^{d}\times^{K})$ of Gaussian probability measures. Therefore, in the particular case where $\mu_{0}$ is Gaussian, which is a standing assumption in this paper, and the operators $\Psi$ and $h$ are linear, the mean field ensemble Kalman filter (2.5) reproduces the exact filtering distribution (1.9), which is then the Kalman filter itself; indeed this is what we show in Subsection 2.1. In the next section we provide error bounds between (1.9) and (2.5) when $\Psi$ and $h$ are not assumed to be linear.
Our theorems in Subsections 2.3 and 2.4 concern relationships between the true filter (1.9) and the mean-field ensemble Kalman filter (2.5). We study the setting in which $\Psi$ and $h$ are not assumed to be linear, so that the true filter is not Gaussian, in Subsection 2.3; and then we study small perturbations away from the Gaussian setting that arise when the vector fields $\Psi$ and $h$ are close to constant in Subsection 2.4. We recall that that $B_{L^{\infty}}(m,r)$ denotes the $L^{\infty}$ ball of radius $r$, centred at $m$. The theorems hold under the following set of assumptions.
Assumption H.
The following assumptions hold on the data $\{y_{j}^{\dagger}\}$, the vector fields $(\Psi,h)$ and the covariances $(\mathit{\Sigma},\mathit{\Gamma}).$
  1. (H1)
    Fix positive integer $J.$ The data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ lies in set $B_{y}\subset^{KJ}$ defined, for some $\kappa_{y}>0$, by
    $B_{y}:=\left\{Y^{\dagger}\in^{KJ}:\max_{j\in}\lvert y^{\dagger}_{j}\rvert\leqslant\kappa_{y}\right\}.$
  2. (H2)
    The function $\Psi$ satisfies $\Psi(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\in B_{\Psi}:=B_{L^{\infty}}(0,\kappa_{\Psi})$ for some $\kappa_{\Psi}>0$.
  3. (H3)
    The function $h$ satisfies $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\in B_{h}:=B_{L^{\infty}}(0,\kappa_{h})$ for some $\kappa_{h}>0$.
  4. (H4)
    The covariance matrices $\mathit{\Sigma}$ and $\mathit{\Gamma}$ appearing in (2.2) are positive definite: $\mathit{\Sigma}\succcurlyeq\sigma\mathit{I}_{d}$ and $\mathit{\Gamma}\succcurlyeq\gamma\mathit{I}_{K}$ for positive $\sigma$ and $\gamma$.
2.3 Stability Theorem: Ensemble Kalman Filter
Roughly speaking, our main result states that if the true filtering distributions $(\mu_{j})_{j\in}$ are close to Gaussian after appropriate lifting to the state/data space then the distribution $\mu^{K}_{j}$ given by the mean field ensemble Kalman filter (2.5) is close to the true filtering distribution $\mu_{j}$ given by (1.9) for all $j\in\llbracket 0,J\rrbracket$. Recall that $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ denotes the $C^{0,1}$ semi-norm, namely the Lipschitz constant.
Theorem 4 (Stability for the Mean Field Ensemble Kalman Filter).
Assume that the probability measures $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ are obtained respectively from the dynamical systems (1.9) and (2.5), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\qquad\mu^{K}_{j+1}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{K}.$
If Assumption H holds and $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$, then there exists $C=C(J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\mathit{\Sigma},\mathit{\Gamma})$ such that
$d_{g}(\mu^{K}_{J},\mu_{J})\leqslant C\max_{j\in}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
Remark 5.
The stability theorem shows that the error made by the ensemble Kalman filter is controlled by the difference between the true filter, lifted to the joint space of state and observations, and its Gaussian projection. The constant $C$ in the stability estimate is uniform across realizations of the data from $B_{y}.$ Indeed, since $h$ is assumed bounded, the bound $\kappa_{y}$ on the data will hold with probability exponentially close to $1$, [^1] With respect to the noise realization leading to the data. for $\kappa_{y}\gg 1$, when there is no model mis-specification. Relaxing the assumptions of bounded $\Psi$ and $h$, for example to the setting of linear plus bounded functions, is technically challenging and will be left for future work. Likewise relaxing positivity assumptions on the noise leads to substantial technical obstacles, deferred for future work. And finally, relaxing the assumption that $J$ is finite will require further structural assumptions on the long-time behaviour of the filter, and is also deferred for future work.
The proof presented hereafter relies on a number of auxiliary results, which are summarized below and proved in Appendix B.
  1. 1.
    For any probability measure $\mu$, the first moments of the probability measures $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their second moments are bounded both from above and from below. The constants in these bounds depend only on the parameters $\kappa_{\Psi}$, $\kappa_{h}$, $\mathit{\Sigma}$ and $\mathit{\Gamma}$. See Lemmas 15 and 16.
  2. 2.
    For any Gaussian measure $\mu\in\mathcal{G}(^{d}\times^{K})$, it holds that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma 19 and also [8].
  3. 3.
    The map $\mathsf{P}$ is globally Lipschitz on $\mathcal{P}(^{d})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{P}}$ depending only on the parameters $\kappa_{\Psi}$ and $\mathit{\Sigma}$. See Lemma 21.
  4. 4.
    The map $\mathsf{Q}$ is globally Lipschitz on $\mathcal{P}(^{d})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{Q}}$ depending only on the parameters $\kappa_{h}$ and $\mathit{\Gamma}$. See Lemma 22.
  5. 5.
    The map $\mathsf{B}_{j}$ satisfies for any $\pi\in\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}(^{d}\times^{K})$ the bound
    $\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{B}_{j}\mathsf{G}\pi,\mathsf{B}_{j}\pi)\leqslant C_{\mathsf{B}}d_{g}(\mathsf{G}\pi,\pi),$
    where $C_{\mathsf{B}}=C_{\mathsf{B}}(\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$. This statement concerns the stability of the $\mathsf{B}_{j}$ operator between a measure and its Gaussian approximation. See Lemma 23.
  6. 6.
    The map $\mathsf{T}_{j}$ satisfies the following bound: for all $R\geqslant 1$, it holds for all probability measures $\pi\in\mathcal{P}_{R}(^{d}\times^{K})$ and $p\in\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}(^{d}\times^{K})$ that
    $\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p),$
    for a constant $L_{\mathsf{T}}=L_{\mathsf{T}}(R,\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$. This statement may be viewed as a local Lipschitz continuity result in the case where the second argument of $d_{g}$ is restricted to the range of $\mathsf{Q}\mathsf{P}$. See Lemma 24.
Proof of Theorem 4.
In what follows we refer to the preceding itemized list to clarify the proof. For notational simplicity it is helpful to define the following measure of the difference between the true filtering distribution and its Gaussian projection:
$\varepsilon:=\max_{j\in}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$ (2.6)
Assume throughout the following that $j\in.$ The main idea of the proof results from the following use of the triangle inequality
$d_{g}(\mu^{K}_{j+1},\mu_{j+1})=d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{K},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$ (2.7a)
$\quad\leqslant d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{K},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)+d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)+d_{g}\left(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right).$ (2.7b)
We have used the fact that the second argument of the second term on the right-hand side indeed coincides with the first argument of the third term because $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by item 2 (Lemma 19.) By item 1 (Lemma 16), there is a constant $R\geqslant 1$, depending on the covariance matrices $\mathit{\Sigma}$, $\mathit{\Gamma}$ and the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, such that $\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}_{R}(^{d}\times^{K})$. By items 3, 4 and 6 (Lemmas 24, 22 and 21), the first term in (2.7b) satisfies
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{K},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$ $\leqslant L_{\mathsf{T}}(R)L_{\mathsf{Q}}L_{\mathsf{P}}d_{g}\left(\mu_{j}^{K},\mu_{j}\right),$ (2.8)
where, for conciseness, we omitted the dependence of the constants on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma}$. Equation (2.8) establishes that the composition of maps $\mathsf{T}_{j}\mathsf{Q}\mathsf{P}$ is globally Lipschitz over $\mathcal{P}(^{d})$. Since $\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\in\mathcal{G}_{R}(^{d}\times^{K})$ by definition of $R$, the second term in (2.7b) may be bounded using item 6 (Lemma 24) and the definition in (2.6) of $\varepsilon$:
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}(R)d_{g}\left(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}(R)\varepsilon.$
Finally, the third term in (2.7b) can be bounded using item 5 (Lemma 23) and the definition in (2.6) of $\varepsilon$:
$d_{g}\left(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant C_{\mathsf{B}}d_{g}\left(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant C_{\mathsf{B}}\varepsilon.$
Therefore, letting $\ell=L_{\mathsf{T}}(R)L_{\mathsf{Q}}L_{\mathsf{P}}$, we have shown that
$d_{g}(\mu^{K}_{j+1},\mu_{j+1})$ $\leqslant\ell d_{g}(\mu^{K}_{j},\mu_{j})+\bigl{(}L_{\mathsf{T}}(R)+C_{\mathsf{B}}\bigr{)}\varepsilon$
and the conclusion follows from the discrete Gronwall lemma, since $\mu_{0}=\mu_{0}^{K}.$
2.4 Approximation Theorem: Ensemble Kalman Filter
Theorem 4 shows that the ensemble Kalman filter error can be made arbitarily small if the true filter is arbitarily close to its Gaussian projection, in state-observation space. This ’closeness to Gaussian’ assumption can be satisfied in our setting of bounded vector fields by considering small perturbations of constant vector fields and is the content of the following corollary. The result provides a first step in the analysis of the accuracy of the ensemble Kalman filter, as an approximation of the true filter; desirable generalizations of what we prove here are discussed in the conclusions, in Section 4.
Corollary 6 (Accuracy for the Mean Field Ensemble Kalman Filter).
Let $(\mathit{\Sigma},\mathit{\Gamma})$ satisfy Assumption (H4). Suppose that $\Psi_{0}\colon^{d}\to^{d}$ and $h_{0}\colon^{d}\to^{K}$ are functions taking constant values and denote by $B_{\Psi_{0},h_{0}}(r)$ the set of all functions $(\Psi,h)$ satisfying $\Psi\in B_{L^{\infty}}(\Psi_{0},r)$, $h\in B_{L^{\infty}}(h_{0},r)$ and Assumptions (H2) and (H3). Assume also that $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$ and denote by $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ the probability measures obtained respectively from the dynamical systems (1.9) and (2.5), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\qquad\mu^{K}_{j+1}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{K}.$
Then for any $\epsilon>0$ there exists $\delta>0$ such that
$\sup_{Y^{\dagger}\in B_{y}}\sup_{(\Psi,h)\in B_{\Psi_{0},h_{0}}(\delta)}d_{g}(\mu^{K}_{J},\mu_{J})\leqslant\epsilon.$
Proof.
The result follows from Theorem 4 and Proposition 31 in Appendix C. ∎
3 Generalization: Gaussian Projected Filter
We generalize the main result to the Gaussian projected filter defined in [8]. This algorithm may be viewed as a mean field version of the unscented Kalman filter [37]. Using a similar approach to that adopted in the previous section, we prove a similar stability theorem and accuracy corollary. The Gaussian projected filter is defined by the iteration
$\mu^{G}_{j+1}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu^{G}_{j}.$ (3.1)
This iteration is obtained from (1.9) by inserting a projection onto Gaussians before the conditioning step $\mathsf{B}_{j}.$ Because conditioning of Gaussians on linear observations preserves Gaussianity, the preceding map generates sequence of measures $\{\mu^{G}_{j}\}$ in $\mathcal{G}(^{d}).$
Remark 7.
As shown in Lemma 20, the evolution (3.1) may also be rewritten in the form
$\mu^{G}_{j+1}=\mathsf{G}\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu^{G}_{j}.$ (3.2)
This shows that the Gaussian projected filter may also be obtained from the mean field ensemble Kalman filter (2.5) by adding a Gaussian projection step after the conditioning.
Like (2.5) the filter (3.1) reproduces the true filtering distributions when $\Psi$ and $h$ are linear, assuming that $\mu_{0}$ is Gaussian. The following theorem quantifies closeness of the Gaussian projected filter to the true filtering distribution when the linearity assumption on $\Psi$ and $h$ is relaxed.
Theorem 8 (Stability for the Gaussian Projected Filter).
Assume that the probability measures $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ are obtained respectively from the dynamical systems (1.9) and (3.1), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{G}$. That is,
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\qquad\mu^{G}_{j+1}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}^{G}.$
If Assumption H holds, then there exists $C=C(J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$ such that
$d_{g}(\mu^{G}_{J},\mu_{J})\leqslant C\max_{j\in}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
The proof of this error estimate relies on the auxiliary results Lemmas 16, 21, 22 and 23 already discussed, as well as the following two additional results.
  • The map $\mathsf{G}$ is locally Lipschitz for the metric $d_{g}$, in the sense that for any $R\geqslant 1$, this map is Lipschitz continuous over the set $\mathcal{P}_{R}(^{d})$ given in (1.1). The associated Lipschitz constant is denoted by $L_{\mathsf{G}}=L_{\mathsf{G}}(R)$ and diverges as $R\to\infty$. This result is proved in Lemma 26, which relies on an auxiliary result shown in Lemma 25 on the distance between Gaussians in the $d_{g}$ metric.
  • The map $\mathsf{B}_{j}$ is locally Lipschitz for the metric $d_{g}$ over Gaussians, in the sense that for any $R\geqslant 1$ and any $j\in$, this map is Lipschitz continuous over $\mathcal{G}_{R}(^{d}\times^{K})$. The associated Lipschitz constant is denoted by $L_{\mathsf{B}}=L_{\mathsf{B}}(R,\kappa_{y})$. See Lemma 29.
Proof.
We obtain by the triangle inequality that
$d_{g}(\mu_{j+1}^{G},\mu_{j+1})$ $=d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}^{G},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j})$
$\leqslant d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}^{G},\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j})+d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}).$
It follows from Lemmas 16, 21, 22, 26 and 29 that the composition of maps $\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}$ is globally Lipschitz continuous on $\mathcal{P}(^{d})$ with a constant $\ell=\ell(R,\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$, where $R=R(\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$ is a positive constant such that $\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}_{R}(^{d}\times^{K})$. Therefore, the first term on the right hand side may be bounded by
$d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}^{G},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j})\leqslant\ell d_{g}(\mu_{j}^{G},\mu_{j}),$
For notational simplicity we again let $\varepsilon$ be as in (2.6). Using Lemma 23 and definition of $\varepsilon$, the second term may be bounded from above by
$d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j})$ $\leqslant C_{\mathsf{B}}\,d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j})\leqslant C_{\mathsf{B}}\,\varepsilon,$
where $C_{\mathsf{B}}=C_{\mathsf{B}}(\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$ is the constant from Lemma 23. The proof can then be concluded in the same way as that of Theorem 4. ∎
Corollary 9 (Accuracy for the Gaussian Projected Filter).
Let $(\mathit{\Sigma},\mathit{\Gamma})$ satisfy Assumption (H4). Suppose that $\Psi_{0}\colon^{d}\to^{d}$ and $h_{0}\colon^{d}\to^{K}$ are functions taking constant values and denote by $B_{\Psi_{0},h_{0}}(r)$ the set of all functions $(\Psi,h)$ satisfying $\Psi\in B_{L^{\infty}}(\Psi_{0},r)$, $h\in B_{L^{\infty}}(h_{0},r)$ and Assumptions (H2) and (H3). Assume that the probability measures $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ are obtained respectively from the dynamical systems (1.9) and (3.1), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\qquad\mu^{G}_{j+1}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}^{G}.$
Then for any $\epsilon>0$ there exists $\delta>0$ such that
$\sup_{Y^{\dagger}\in B_{y}}\sup_{(\Psi,h)\in B_{\Psi_{0},h_{0}}(\delta)}d_{g}(\mu^{G}_{J},\mu_{J})\leqslant\epsilon.$
4 Discussion and Future Directions
We have provided the first analysis of the error incurred by ensemble Kalman filters, as approximations of true filtering distribution, beyond the linear Gaussian setting. We have employed an appropriate weighted TV metric and obtained new stability estimates in this metric, in order to establish the approximation results. Our framing of the problem is motivated by the framing of the analysis of the particle filter contained in [50, Section 1], a systematization of the original proof of convergence of the particle filter appearing in [14]. Although it introduces new methodology and theoretical results for nonlinear Kalman filters, our work leaves open numerous avenues for further analysis; we now highlight those that we identify as particularly important. These remaining open problems are substantial, but the framework we map out in this paper is an appropriate one in which to address them.
  1. (i)
    Our theorems concern the mean field limit of the ensemble Kalman filter; it is of interest to study finite particle approximations of the mean field limit, along the lines of the work in [44, 6] and analogous continuous time analyses in [17, 18, 19]. Analysis of mean field limits of interacting particle systems is an established field; interfacing the natural metrics employed for such analyses (Wasserstein) with those employed here (weighted TV) will be required.
  2. (ii)
    We have made strong assumptions on $\Phi(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ and $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ in this paper; developing proofs which relax these assumptions, is important. It will likely require use of Lyapunov functions and lead to bounds which hold in high probability with respect to the observed data.
  3. (iii)
    Our error bounds are for a finite number of steps and, as is typical of finite time error estimates that employ a consistency plus stability implies convergence approach, lead to error constants which grow exponentially with the time horizon. Generalizing to long-time estimates, using assumptions on the long-term stability, or even ergodicity, of the true filter, would be of great interest. Related issues are widely studied for the true filtering distribution; see [49, 15, 62, 61, 10, 53] for example. In the context of the EnKF they are investigated in the paper [13] which identifies linear and Gaussian filtering problems in which it is possible to generalize the large particle asymptotic analyses of [44, 46]. There is also analysis of the EnKF for nonlinear non-Gaussian problems, such as the data assimilation problem for the Navier-Stokes equation, but this work concerns only accuracy of state estimation, not the entire filtering distribution [39].
  4. (iv)
    We have assumed a model for evolution of the state which employs additive Gaussian noise. Generalizing to a general Markov chain would be valuable. Relatedly it is of interest to relax the assumption of including noise in the dynamical system for the state, to allow for deterministic dynamics.
  5. (v)
    We have studied a (widely used) version of the ensemble Kalman filter which employs a specific transport map to approximate the conditioning step in the filter. Other transport maps are also used in practice, such as that leading to the ensemble square root filter; see as outlined in [8, Section 2]. Analyzing these other methods would be of great interest.
  6. (vi)
    Filtering can be used to solve inverse problems, as outlined in [8, Section 4]; in particular some of these methods rely on filtering over the infinite time horizon, the analysis of which will require new ideas such as in [21].
  7. (vii)
    Continuous time versions of our analysis would be of interest, and their relationship to the Kushner–Stratonovich equation [11]; the paper [13], which studies mean field limits of the ensemble Kalman–Bucy filter, may be important in this context; furthermore study of the ensemble Kalman sampler [27, 28, 9] for inverse problems, would be of interest.
  8. (viii)
    The papers [44] and [19] propose reweighting of ensemble Kalman methods to make them unbiased; further analysis of this idea, and the development of new methods that can carry this out in a derivative-free fashion, would be of interest.
  9. (ix)
    Ensemble Kalman filters in practice employ very small numbers of ensemble members, and use covariance (spatial) localization when the unknown states are fields; developing analyses which account for low rank approximations and localization would be a valuable step for the field. Furthermore it would also be of interest to study the problem of quantization of measures [31] in the context of filtering; however this is a very hard problem even in simple low dimensional settings [7] and studying it for the evolution of the filtering distribution will require substantial new ideas.
  10. (x)
    Particle filters suffer from a curse of dimensionality on certain families of high-dimensional problems [4, 56]. Studying ensemble Kalman methods, to determine whether they ameliorate this issue in the near-Gaussian setting, would be of interest. There are a variety of different models that can be employed to characterize families of high dimensional filtering problems [52] and such structural assumptions will undoubtedly affect the results that might be obtained for ensemble Kalman methods.
  11. (xi)
    Exploiting small noise or large data limits, to establish that the error (2.6) is small, and then using this fact to analyze the error in the ensemble Kalman filter, would be of great interest.
Acknowledgments
JAC was supported by the Advanced Grant Nonlocal-CPD (Nonlocal PDEs for Complex Particle Dynamics: Phase Transitions, Patterns and Synchronization) of the European Research Council Executive Agency (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 883363). JAC was also partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/T022132/1 and EP/V051121/1. FH is supported by start-up funds at the California Institute of Technology. FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM. The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship. UV is partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grant ANR-21-CE40-0006 (SINEQ).
Appendix A Auxiliary Results
We begin by presenting an elementary result used throughout the article.
Lemma 10.
Suppose that $X$ is a random variable with values in ${}^{d}$ and finite second moment, and let $m:=\mathbf{E}[X]$. Then
$\mathbf{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}=\mathbf{E}\bigl{[}XX^{\mathsf{T}}\bigr{]}-mm^{\mathsf{T}}$ (A.1)
and
$\forall a\in^{d},\qquad\mathbf{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}\succcurlyeq\mathbf{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}.$ (A.2)
Proof.
We have
$\mathbf{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}$ $=\mathbf{E}\Bigl{[}\bigl{(}(X-m)+(m-a)\bigr{)}\bigl{(}(X-m)+(m-a)\bigr{)}^{\mathsf{T}}\Bigr{]}$
$=\mathbf{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}+(m-a)(m-a)^{\mathsf{T}}.$
Taking $a=0$, we obtain (A.1). In addition, since the second term in the last expression is positive semidefinite, the inequality (A.2) follows. ∎
The following lemma is very similar to [2, Lemma 3.1]; the only difference is that the weight on the left-hand side is given by $1+\lvert u\rvert^{2}$ instead of $u^{2}$. We give the proof for completeness.
Lemma 11 (Generalized Pinsker inequality).
Let $g(u)=1+\lvert u\rvert^{2}$ and assume that $\mu_{1}$, $\mu_{2}$ are probability measures over ${}^{d}$ satisfying $\mu_{1}[g^{2}]<\infty$ and $\mu_{2}[g^{2}]<\infty$. Then, if $\mu_{1}\ll\mu_{2}$,
$d_{g}(\mu_{1},\mu_{2})^{2}\leqslant 2\bigl{(}\mu_{1}\left[g^{2}\right]+\mu_{2}\left[g^{2}\right]\bigr{)}\mathrm{KL}(\mu_{1}\|\mu_{2}).$
Proof.
Denote the density by $f:=\frac{\mathrm{d}\mu_{1}}{\mathrm{d}\mu_{2}}$, noting that this is non-negative. Applying Taylor’s formula to the function $\ell\colon u\mapsto u\log u$ around $u=1$, we deduce that
$\forall u\geqslant 0,\qquad u\log u\geqslant(u-1)+\frac{1}{2}\left(\min_{v\in I}\ell^{\prime\prime}(v)\right)(u-1)^{2}=(u-1)+\frac{(u-1)^{2}}{2\max\{1,u\}},$
where $I=[\min\{1,u\},\max\{1,u\}]$. Therefore
$\mathrm{KL}(\mu_{1}\|\mu_{2})=\int_{{}^{d}}(f\,\log f)(u)\,\mu_{2}(\mathrm{d}u)\geqslant\int_{{}^{d}}(f(u)-1)\,\mu_{2}(\mathrm{d}u)+\frac{1}{2}\int_{{}^{d}}\frac{\lvert f(u)-1\rvert^{2}}{\max\{1,f(u)\}}\,\mu_{2}(\mathrm{d}u),$
Let $\theta(u)=\max\{1,f(u)\}$. The first term on the right-hand side is zero, and so we have that
$d_{g}(\mu_{1},\mu_{2})^{2}$ $=\left(\int_{{}^{d}}g(u)\left\lvert f(u)-1\right\rvert\,\mu_{2}(\mathrm{d}u)\right)^{2}\leqslant\int_{{}^{d}}\lvert g(u)\rvert^{2}\theta(u)\,\mu_{2}(\mathrm{d}u)\,\int_{{}^{d}}\frac{\left\lvert f(u)-1\right\rvert^{2}}{\theta(u)}\,\mu_{2}(\mathrm{d}u)$
[…]
which concludes the proof. ∎
Lemma 12.
Denote by $g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m,S)$ the Lebesgue density of $\mathcal{N}(m,\mathit{S})$ and by $\mathcal{S}^{K}_{\alpha}$ the set of symmetric $K\times K$ matrices $\mathit{M}$ satisfying
$\frac{1}{\alpha}\mathit{I}_{K}\preccurlyeq\mathit{M}\preccurlyeq\alpha\mathit{I}_{K}.$ (A.3)
Then for all $\alpha\geqslant 1$, there exists $L_{\alpha}>0$ such that for all parameters $(c_{1},m_{1},\mathit{S}_{1})\in\times^{K}\times\mathcal{S}^{K}_{\alpha}$ and $(c_{2},m_{2},\mathit{S}_{2})\in\times^{K}\times\mathcal{S}^{K}_{\alpha}$,
$\lVert h\rVert_{\infty}\leqslant L_{\alpha}\lVert h\rVert_{1},\qquad h(y)=c_{1}g(y;m_{1},\mathit{S}_{1})-c_{2}g(y;m_{2},\mathit{S}_{2}).$ (A.4)
Remark 13.
When $c_{2}=0$ then equation (A.4) may be viewed as an inverse inequality. It then simply states that the $L^{\infty}$ norm of the density of a normal random variable is bounded from above by the $L^{1}$ norm, uniformly for all densities from the set of Gaussians with a covariance matrix satisfying (A.3).
Proof.
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context and may change from line to line. The statement is obvious if $c_{1}c_{2}=0$, so we assume from now on that $c_{2}>0$. We divide the proof into three steps.
Step 1: simplification
Since both sides of the inequality (A.4) are invariant under translation, it is sufficient to consider the case where $m_{2}=0$. Note also that
$c_{1}g(y;m_{1},\mathit{S}_{1})-c_{2}g(y;0,\mathit{S}_{2})$
$\qquad=\frac{c_{2}}{\sqrt{\det\mathit{S}_{2}}}\left(\frac{c_{1}}{c_{2}}g\left(\sqrt{\mathit{S}_{2}^{-1}}y;\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}}\right)-g\left(\sqrt{\mathit{S}_{2}^{-1}}y;0,\mathit{I}_{K}\right)\right),$
and so we can also assume without loss of generality that $c_{2}=1$ and $\mathit{S}_{2}=\mathit{I}_{K}$. Indeed, assume that the inequality (A.4) is satisfied for this particular case with a constant $\widetilde{L}_{\alpha}$. Since
$\frac{1}{\alpha^{2}}\mathit{I}_{K}\preccurlyeq\frac{1}{\alpha\lVert\mathit{S}_{2}\rVert}\mathit{I}_{K}\preccurlyeq\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}}\preccurlyeq\alpha\lVert\mathit{S}_{2}^{-1}\rVert\mathit{I}_{K}\preccurlyeq\alpha^{2}\mathit{I}_{K},$
we deduce that if $(\mathit{S}_{1},\mathit{S}_{2})\in\mathcal{S}^{K}_{\alpha}\times S^{K}_{\alpha}$ for some $\alpha>0$, then
$\left(\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}},\mathit{I}_{K}\right)\in\mathcal{S}^{K}_{\alpha^{2}}\times\mathcal{S}^{K}_{\alpha^{2}}.$
Therefore, using the change of variable $y\mapsto\sqrt{\mathit{S}_{2}^{-1}}y$ together with (A.4) in the aforementioned particular case, we obtain that
$\lVert c_{1}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-c_{2}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\rVert_{\infty}$
$\qquad=\frac{\lvert c_{2}\rvert}{\sqrt{\det\mathit{S}_{2}}}\left\lVert\frac{c_{1}}{c_{2}}g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}}\right)-g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K}\right)\right\rVert_{\infty}$
$\qquad\leqslant\frac{\widetilde{L}_{\alpha^{2}}\lvert c_{2}\rvert}{\sqrt{\det\mathit{S}_{2}}}\left\lVert\frac{c_{1}}{c_{2}}g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{S_{2}^{-1}}S_{1}\sqrt{S_{2}^{-1}}\right)-g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K}\right)\right\rVert_{1}$
$\qquad=\frac{\widetilde{L}_{\alpha^{2}}}{\sqrt{\det\mathit{S}_{2}}}\left\lVert c_{1}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-c_{2}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\right\rVert_{1}\leqslant\alpha^{K/2}\widetilde{L}_{\alpha^{2}}\left\lVert c_{1}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-c_{2}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\right\rVert_{1},$
and so the bound (A.4) is valid in general with $L_{\alpha}=\alpha^{K/2}\widetilde{L}_{\alpha^{2}}$.
Step 2: contradiction hypothesis
For simplicity, we use from now on the notations $c,m,\mathit{S}$ in place of $c_{1},m_{1},\mathit{S}_{1}$. It remains to show that there exists for all $\alpha>0$ a constant $\widetilde{L}_{\alpha}$ such that
$\lVert h\rVert_{\infty}\leqslant\widetilde{L}_{\alpha}\lVert h\rVert_{1},\qquad h(y)=cg(y;m,\mathit{S})-g(y;0,\mathit{I}_{K}).$
for all $(c,m,\mathit{S})\in\times^{K}\times\mathcal{S}^{K}_{\alpha}$. Assume for contradiction that there is a sequence of parameters $(c_{n},m_{n},\mathit{S}_{n})_{n\geqslant 0}$ in ×$\times^{K}\times\mathcal{S}^{K}_{\alpha}$ and not containing $(1,0,\mathit{I}_{K})$ and such that
$\frac{\lVert h_{n}\rVert_{\infty}}{\lVert h_{n}\rVert_{1}}\xrightarrow[n\to\infty]{}\infty,\qquad h_{n}(y):=c_{n}g(y;m_{n},\mathit{S}_{n})-g(y;0,\mathit{I}_{K}).$ (A.5)
Since $\lVert h_{n}\rVert_{\infty}\leqslant c_{n}/\sqrt{(2\pi)^{K}\det\mathit{S}_{n}}+1/\sqrt{(2\pi)^{K}}$ and
$\lvert c_{n}-1\rvert=\left\lvert\int_{{}^{K}}h_{n}\,\mathrm{d}y\right\rvert\leqslant\lVert h_{n}\rVert_{1}$
we have
$\frac{\lVert h_{n}\rVert_{\infty}}{\lVert h_{n}\rVert_{1}}\leqslant\frac{c_{n}/\sqrt{\det\mathit{S}_{n}}+1}{\sqrt{(2\pi)^{K}}\lvert c_{n}-1\rvert}\leqslant\frac{c_{n}\sqrt{\alpha^{K}}+1}{\sqrt{(2\pi)^{K}}\lvert c_{n}-1\rvert}.$
The right-hand side can diverge only if $c_{n}\to 1$ in the limit $n\to\infty$, and so this must necessarily hold. If $c_{n}\to 1$ then $\lVert h_{n}\rVert_{\infty}$ is uniformly bounded from above as $n\to\infty$ implying by (A.5) that $\lVert h_{n}\rVert_{1}\to 0$. Now notice that
$h_{n}=\Bigl{(}c_{n}g(y;m_{n},\mathit{S}_{n})-g(y;m_{n},\mathit{S}_{n})\Bigr{)}+\Bigl{(}g(y;m_{n},\mathit{S}_{n})-g(y;0,\mathit{I}_{K})\Bigr{)},$
The $L^{1}(^{K})$ norm of the first term on the right-hand side tends to 0 as $n\to\infty$, so that of the second term must also converge to 0 in $L^{1}(^{K})$. This term may be bounded in norm by using [16, Proposition 2.2], which gives a lower bound on the total variation distance between Gaussians. In our setting, this result implies
$\frac{1}{2}\lVert g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{n},\mathit{S}_{n})-g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K})\rVert_{1}$ $\geqslant 1-\frac{\det\mathit{S}_{n}^{1/4}\det\mathit{I}_{K}^{1/4}}{\det\left(\frac{\mathit{S}_{n}+\mathit{I}_{K}}{2}\right)^{1/2}}\exp\left(-\frac{1}{8}m_{n}^{\mathsf{T}}\left(\frac{\mathit{S}_{n}+\mathit{I}_{K}}{2}\right)^{-1}m_{n}\right).$ (A.6)
Since the left-hand side of this inequality converges to 0, it necessarily holds that $m_{n}\to 0$ and $\mathit{S}_{n}\to\mathit{I}_{K}$. Indeed, the right-hand side is bounded from below by
$1-\frac{\det\mathit{S}_{n}^{1/4}\det\mathit{I}_{K}^{1/4}}{\det\left(\frac{\mathit{S}_{n}+\mathit{I}_{K}}{2}\right)^{1/2}}=1-\prod_{i=1}^{K}\frac{\lambda_{n,i}^{1/4}}{\left(\frac{\lambda_{n,i}+1}{2}\right)^{1/2}}=1-\prod_{i=1}^{K}\sqrt{\frac{\sqrt{\lambda_{n,i}}}{\frac{\lambda_{n,i}+1}{2}}},$
where $(\lambda_{n,i})_{i=1}^{K}$ are the eigenvalues of $\mathit{S}_{n}$. All the factors in the product are bounded from above by 1 by the arithmetic mean-geometric mean inequality. Since the left-hand side of (A.6) converges to 0, all these factors must converge to 1 in the limit as $n\to\infty$, otherwise the right-hand side of (A.6) would be bounded away from 0 uniformly in $n$. Thus $\lambda_{n,i}\to 1$ for all $i\in\llbracket 1,K\rrbracket$, and so $S_{n}$ converges to $\mathit{I}_{K}$ as $n\to\infty$. This implies that the factor multiplying the exponential in (A.6) converges to 1, so the exponential itself must converge to 1, whence we deduce $m_{n}\to 0$.
Step 3: reaching a contradiction
Since $(c_{n},m_{n},\mathit{S}_{n})$ converges to $(1,0,\mathit{I}_{K})$ as $n\to\infty$, a contradiction can be reached by showing that $\lVert h\rVert_{\infty}/\lVert h\rVert_{1}$ is bounded from above uniformly for any $(c,m,\mathit{S})\neq(1,0,\mathit{I}_{K})$ in a small neighborhood of $(1,0,\mathit{I}_{K})$. To this end, we introduce the following notation for any $(c,m,\mathit{S})\in\times^{K}\times\mathcal{S}^{K}_{\alpha}$:
$\upsilon_{0}=\frac{c}{\sqrt{\det\mathit{S}}}-1,\qquad\upsilon_{1}=m,\qquad\mathit{\Upsilon}_{2}=\frac{\mathit{I}_{K}-\mathit{S}^{-1}}{2},\qquad r=\max\Bigl{\{}\lvert\upsilon_{0}\rvert,\lvert\upsilon_{1}\rvert,\lVert\mathit{\Upsilon}_{2}\rVert\Bigr{\}}.$ (A.7)
Although we omit the arguments $(c,m,\mathit{S})$ for conciseness, these quantities are viewed as functions over ×$\times^{K}\times\mathcal{S}^{K}_{\alpha}$. Then
$h(y)$ $=\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\left((1+\upsilon_{0})\exp\left(-\frac{1}{2}(y-\upsilon_{1})^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})(y-\upsilon_{1})+\frac{y^{\mathsf{T}}y}{2}\right)-1\right).$ (A.8)
The argument of the exponential within the brackets can be rewritten as
$a=a(y;\upsilon_{1},\mathit{\Upsilon}_{2}):=\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y-2\upsilon_{1}^{\mathsf{T}}\mathit{\Upsilon}_{2}y-\frac{1}{2}\upsilon_{1}^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})\upsilon_{1},$
which may be bounded from above as
$\left\lvert a(y;\upsilon_{1},\mathit{\Upsilon}_{2})\right\rvert$ $\leqslant\left(r+r+2r^{2}+\frac{1}{2}r^{2}+r^{3}\right)\bigl{(}1+\lvert y\rvert^{2}\bigr{)},$ (A.9a)
$\left\lvert\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert$ $\leqslant 2r\bigl{(}1+\lvert y\rvert^{2}\bigr{)},$ (A.9b)
$\left\lvert 2\upsilon_{1}^{\mathsf{T}}\mathit{\Upsilon}_{2}y+\frac{1}{2}\upsilon_{1}^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})\upsilon_{1}\right\rvert$ $\leqslant\left(2r^{2}+\frac{1}{2}r^{2}+r^{3}\right)\bigl{(}1+\lvert y\rvert^{2}\bigr{)}$ (A.9c)
Let $\mathcal{A}_{\varepsilon}$ denote the neighborhood $\bigl{\{}(c,m,\mathit{S})\in\times^{K}\times\mathcal{S}^{\alpha}_{K}:r\leqslant\varepsilon\bigr{\}}$. Employing the triangle inequality together with the elementary inequality $\lvert\operatorname{e}^{x}-1-x\rvert\leqslant\frac{x^{2}}{2}\operatorname{e}^{\lvert x\rvert},$ which is valid for any $x\in$ and obtained from a Taylor expansion, we deduce that,
$\forall(c,m,\mathit{S})\in\mathcal{A}_{1},\qquad\forall y\in^{K},\qquad$
$\qquad\left\lvert\operatorname{e}^{a}-1-\upsilon_{1}^{\mathsf{T}}y-y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert$ $\leqslant\left\lvert 2\upsilon_{1}^{\mathsf{T}}\mathit{\Upsilon}_{2}y+\frac{1}{2}\upsilon_{1}^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})\upsilon_{1}\right\rvert+\frac{1}{2}a^{2}\operatorname{e}^{\lvert a\rvert}$
$\leqslant 4r^{2}\left(1+\lvert y\rvert^{2}\right)+\frac{1}{2}(6r)^{2}\bigl{(}1+\lvert y\rvert^{2}\bigr{)}^{2}\exp\biggl{(}6r\left(1+\lvert y\rvert^{2}\right)\biggr{)}$
$\leqslant 22r^{2}\left(1+\lvert y\rvert^{2}\right)^{2}\exp\biggl{(}6r\left(1+\lvert y\rvert^{2}\right)\biggr{)}.$
We have used the fact that, since $(c,m,\mathit{S})\in\mathcal{A}_{1}$, we have $r\leqslant 1.$ From this equation and (A.8) and the expression
$h(y)-\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\left(\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right)$
$\qquad=\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\Bigl{(}\operatorname{e}^{a}-1-\upsilon_{1}^{\mathsf{T}}y-y^{\mathsf{T}}\mathit{\Upsilon}_{2}y+\upsilon_{0}(\operatorname{e}^{a}-1)\bigr{)}\Bigr{)},$
we deduce that there exists $\varepsilon_{1}\in(0,1)$ sufficiently small such that for all $(c,m,\mathit{S})\in\mathcal{A}_{\varepsilon_{1}}$, the following inequality holds:
$\forall y\in^{K},\qquad\left\lvert h(y)-\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\left(\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right)\right\rvert\leqslant Cr^{2}\exp\left(-\frac{y^{\mathsf{T}}y}{4}\right).$ (A.10)
Therefore, it holds that $\lVert h\rVert_{\infty}$ is bounded from above by $Cr$ uniformly for $(c,m,\mathit{S})$ in $\mathcal{A}_{\varepsilon_{1}}$. In order to reach a contradiction and conclude the proof, it is therefore sufficient to prove that $\lVert h\rVert_{1}$ is bounded from below by $Cr$ for $(c,m,\mathit{S})$ in $\mathcal{A}_{\varepsilon_{2}}$ for some $\varepsilon_{2}\in(0,1)$. It follows from (A.10) that
$\lVert h\rVert_{1}=\int_{{}^{K}}\left\lvert\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert\,g(y;0,\mathit{I}_{K})\,\mathrm{d}y+\mathcal{O}(r^{2}).$
Let $\mathcal{I}$ denote the integral on the right-hand side. It holds that
$\frac{\mathcal{I}}{r}=\int_{{}^{K}}\left\lvert\left(\frac{\upsilon_{0}}{r}\right)+\left(\frac{\upsilon_{1}}{r}\right)^{\mathsf{T}}y+y^{\mathsf{T}}\left(\frac{\mathit{\Upsilon}_{2}}{r}\right)y\right\rvert\,g(y;0,\mathit{I}_{K})\,\mathrm{d}y.$
By definition of $r$, all of the coefficients in parentheses are bounded from above by $1$, in absolute value, vector norm and matrix norm respectively; furthermore one of the terms is necessarily equal to $1$ by one of these measures of size. The conclusion then follows from the lower bound (A.11), which is proved in the next step.
Step 4: proof of the auxiliary inequality
In this step, we prove that there exists a constant $C>0$ such that the following inequality holds for all $(\upsilon_{0},\upsilon_{1},\mathit{\Upsilon}_{2})\in\times^{K}\times^{K\times K}$ satisfying $\max\bigl{\{}\lvert\upsilon_{0}\rvert,\lvert\upsilon_{1}\rvert,\lVert\mathit{\Upsilon}_{2}\rVert\bigr{\}}=1$:
$\int_{{}^{K}}\left\lvert\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert\,g(y;0,\mathit{I}_{K})\mathrm{d}y\geqslant C.$ (A.11)
To this end, note that set of parameters $(\upsilon_{0},\upsilon_{1},\mathit{\Upsilon}_{2})$ satisfying $\max\bigl{\{}\lvert\upsilon_{0}\rvert,\lvert\upsilon_{1}\rvert,\lVert\mathit{\Upsilon}_{2}\rVert\bigr{\}}=1$ is a compact subset, say $\mathcal{K}$, of ×$\times^{K}\times^{K\times K}$. In addition, the left-hand side of (A.11) is continuous with respect to the parameters. Therefore, the infimum of the integral over $\mathcal{K}$ is attained, and since the integral is strictly positive for any tuple of parameters in $\mathcal{K}$, the infimum is necessarily strictly positive. ∎
Lemma 14.
Let $\mathsf{P}$ and $\mathsf{Q}$ denote the operators on probability measures given respectively in (1.5) and (1.7). Suppose that Assumption H is satisfied and that $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$. Then there is $L=L(\kappa_{\Psi},\kappa_{h},\ell_{h},\mathit{\Sigma},\mathit{\Gamma})$ such that for all $(u_{1},u_{2},y)\in^{d}\times^{d}\times^{K}$ and all $\mu\in\mathcal{P}(^{d})$ the density of $p=\mathsf{Q}\mathsf{P}\mu$ satisfies
$\lvert p(u_{1},y)-p(u_{2},y)\rvert\leqslant L\lvert u_{1}-u_{2}\rvert\exp\left(-\frac{1}{4}\left(\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}+\lvert y\rvert_{\mathit{\Gamma}}^{2}\right)\right).$ (A.12)
Proof.
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context, depends only on $\kappa_{\Psi},\kappa_{h},\ell_{h},\mathit{\Sigma},\mathit{\Gamma}$, and may change from line to line. We begin by proving that for all $(u_{1},u_{2},v)\in^{d}\times^{d}\times^{d}$, it holds that
$\left\lvert\exp\left(-\frac{1}{2}\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)-\exp\left(-\frac{1}{2}\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\right\rvert$
$\hskip 142.26378pt\leqslant C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$ (A.13)
Since the operator $\Psi$ is bounded by Assumption H, there is $C$ such that
$\forall(u,v)\in^{d}\times^{d},\qquad\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\leqslant C\exp\left(-\frac{1}{3}\lvert u\rvert_{\mathit{\Sigma}}^{2}\right),$ (A.14)
because the ratio between the exponential functions on the left and right-hand side converges to 00 in the limit as $\lvert u\rvert\to\infty$. By (A.14) it is clear that there is $C>0$ such (A.13) is satisfied whenever $\lvert u_{2}-u_{1}\rvert\geqslant 1$. We now prove that there is $C>0$ such that (A.13) is satisfied when $\lvert u_{2}-u_{1}\rvert\leqslant 1$. It follows from (A.14) and the elementary inequality
$\forall(a,b)\in\times,\qquad\left\lvert\operatorname{e}^{-a}-\operatorname{e}^{-b}\right\rvert\leqslant\lvert b-a\rvert\operatorname{e}^{-\min\{a,b\}}$ (A.15)
that, for any $(u_{1},u_{2},v)\in^{d}\times^{d}\times^{d}$,
$\left\lvert\exp\left(-\frac{1}{2}\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)-\exp\left(-\frac{1}{2}\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\right\rvert$
$\qquad\leqslant\frac{C}{2}\left\lvert\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}-\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right\rvert\exp\left(-\frac{1}{3}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\left\lvert u_{2}\right\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$ (A.16)
Using that $a^{2}-b^{2}=(a-b)(a+b)$ and then the triangle inequality, we obtain
$\left\lvert\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}-\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right\rvert\leqslant\lvert u_{1}-u_{2}\rvert_{\mathit{\Sigma}}\Bigl{(}\lvert u_{1}\rvert_{\mathit{\Sigma}}+\lvert u_{2}\rvert_{\mathit{\Sigma}}+2\lvert\Psi(v)\rvert_{\mathit{\Sigma}}\Bigr{)}.$ (A.17)
The norm $\lvert\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\rvert_{\mathit{\Sigma}}$ is equivalent to the Euclidean norm, and so the right-hand side of (A.16) is bounded from above when $\lvert u_{2}-u_{1}\rvert\leqslant 1$ by
$C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$
Note that the factor in the exponential is now $1/4$ instead of $1/3$, because we used part of the exponential decay in (A.16) in order to control the growth of the term $\lvert u_{1}\rvert_{\mathit{\Sigma}}+\lvert u_{2}\rvert_{\mathit{\Sigma}}$, which appears on the right-hand side of (A.17). It then follows from (A.13) and the definition (1.5) of $\mathsf{P}$ that the density of $\mathsf{P}\mu$ satisfies the following inequality for all $(u_{1},u_{2})\in^{d}\times^{d}$:
$\bigl{\lvert}\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\bigr{\rvert}\leqslant C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$ (A.18)
The above reasoning is now repeated almost verbatim in order to prove that the following inequality holds for all $(u_{1},u_{2},y)\in^{d}\times^{K}\times^{K}$:
$\Bigl{\lvert}\mathcal{N}\bigl{(}h(u_{1}),\mathit{\Gamma}\bigr{)}(y)-\mathcal{N}\bigl{(}h(u_{2}),\mathit{\Gamma}\bigr{)}(y)\Bigr{\rvert}\leqslant C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\lvert y\rvert_{\mathit{\Gamma}}^{2}\right).$ (A.19)
Indeed, since $h$ is bounded by Assumption H, it holds that
$\forall(u,y)\in^{d}\times^{K},\qquad\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)$ $\leqslant C\exp\left(-\frac{1}{3}\lvert y\rvert_{\mathit{\Gamma}}^{2}\right),$ (A.20)
and so by (A.15) we have
$\left\lvert\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u_{1})\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u_{2})\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)\right\rvert$
$\qquad\leqslant\frac{C}{2}\left\lvert\left\lvert y-h(u_{1})\right\rvert_{\mathit{\Gamma}}^{2}-\left\lvert y-h(u_{2})\right\rvert_{\mathit{\Gamma}}^{2}\right\rvert\exp\left(-\frac{1}{3}\lvert y\rvert_{\mathit{\Gamma}}^{2}\right).$ (A.21)
Using that $a^{2}-b^{2}=(a-b)(a+b)$ and then the triangle inequality, we deduce that
$\left\lvert\left\lvert y-h(u_{1})\right\rvert_{\mathit{\Gamma}}^{2}-\left\lvert y-h(u_{2})\right\rvert_{\mathit{\Gamma}}^{2}\right\rvert$ $\leqslant\left\lvert h(u_{2})-h(u_{1})\right\rvert_{\mathit{\Gamma}}\Bigl{(}\left\lvert h(u_{1})\right\rvert_{\mathit{\Gamma}}+\left\lvert h(u_{2})\right\rvert_{\mathit{\Gamma}}+2\lvert y\rvert_{\mathit{\Gamma}}\Bigr{)}$
$\leqslant\lVert\mathit{\Gamma}^{-\frac{1}{2}}\rVert\ell_{h}\lvert u_{2}-u_{1}\rvert\Bigl{(}\left\lvert h(u_{1})\right\rvert_{\mathit{\Gamma}}+\left\lvert h(u_{2})\right\rvert_{\mathit{\Gamma}}+2\lvert y\rvert_{\mathit{\Gamma}}\Bigr{)}.$ (A.22)
Here $\ell_{h}$ is the Lipschitz constant of $h$; see Assumption H. Trading a bit of the exponential decay in (A.21) in order to control the growth from the term $2\lvert y\rvert_{\mathit{\Gamma}}$ in (A.22), we obtain (A.19).
In order to conclude the proof, using the definition of $\mathsf{P}$, we calculate that
$p(u_{1},y)-p(u_{2},y)$ $=\mathsf{P}\mu(u_{1})\,\mathcal{N}\bigl{(}h(u_{1}),\mathit{\Gamma}\bigr{)}(y)-\mathsf{P}\mu(u_{2})\,\mathcal{N}\bigl{(}h(u_{2}),\mathit{\Gamma}\bigr{)}(y)$
$=\bigl{(}\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\bigr{)}\,\mathcal{N}\bigl{(}h(u_{1}),\mathit{\Gamma}\bigr{)}(y)$
$\quad+\mathsf{P}\mu(u_{2})\Bigl{(}\mathcal{N}\bigl{(}h(u_{1}),\mathit{\Gamma}\bigr{)}-\mathcal{N}\bigl{(}h(u_{2}),\mathit{\Gamma}\bigr{)}(y)\Bigr{)}.$
The first and second terms on the right hand side can be bounded by using (A.18) and (A.19), respectively, leading to (A.12). ∎
Appendix B Technical Results for Theorems 4 and 8
We show moment bounds in Subsection B.1, and we prove that $\mathsf{T}_{j}\mu=\mathsf{B}_{j}\mu$ for any Gaussian probability measure $\mu$ in Subsection B.2. Finally, we prove the stability results used in the proofs of Theorems 4 and 8 in Subsections B.3 and B.4, respectively.
B.1 Moment Bounds
Lemma 15 (Moment bounds).
Let $\mu$ denote a probability measure on ${}^{d}$. Under Assumption H, it holds that
$\bigl{\lvert}\mathcal{M}(\mathsf{P}\mu)\bigr{\rvert}\leqslant\kappa_{\Psi},\qquad\mathit{\Sigma}\preccurlyeq\mathcal{C}(\mathsf{P}\mu)\preccurlyeq\kappa_{\Psi}^{2}\mathit{I}_{d}+\mathit{\Sigma}.$ (B.1)
Proof.
From the definition of $\mathsf{P}$ in (1.5), we have that
$\mathcal{M}(\mathsf{P}\mu)=\int_{{}^{d}}u\,\mathsf{P}\mu(u)\,\mathrm{d}u$ $=\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}\int_{{}^{d}}u\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\,\mu(\mathrm{d}v)\,\mathrm{d}u$
$=\int_{{}^{d}}\Psi(v)\,\mu(\mathrm{d}v),$
where the last equality is obtained by changing the order of integration using Fubini’s theorem. Using the first item in Assumption H, we then deduce the first inequality in (B.1). For the second inequality in (B.1), we first note the following inequality which holds for any $m,v\in^{d}$ by Lemma 10:
$\int_{{}^{d}}(u-m)\otimes(u-m)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\Bigr{)}\,\mathrm{d}u$
$\qquad\succcurlyeq\int_{{}^{d}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\Bigr{)}\,\mathrm{d}u;$ (B.2)
The result follows by using the fact that $\Psi(v)$ is the mean under Gaussian $\mathcal{N}\bigl{(}\Psi(v),\mathit{\Sigma}\bigr{)}.$ Now choose $m$ to be the mean under measure $\mathsf{P}\mu$ and note that, by conditioning on $v$ and using (B.1),
$\mathcal{C}(\mathsf{P}\mu)$ $=\int_{{}^{d}}(u-m)\otimes(u-m)\,\mathsf{P}\mu(u)\,\mathrm{d}u$
$=\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}\left(\int_{{}^{d}}(u-m)\otimes(u-m)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
$\succcurlyeq\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}\left(\int_{{}^{d}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
$=\int_{{}^{d}}\mathit{\Sigma}\,\mu(\mathrm{d}v)=\mathit{\Sigma},$
and so $\mathcal{C}(\mathsf{P}\mu)\succcurlyeq\mathit{\Sigma}$. On the other hand, using Lemma 10 again together with the fact that, by the Cauchy-Schwarz inequality, $aa^{\mathsf{T}}\preccurlyeq(a^{\mathsf{T}}a)\mathit{I}_{d}$ for any vector $a\in^{d}$, we have
$\mathcal{C}(\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d}}u\otimes u\,\mathsf{P}\mu(u)\,\mathrm{d}u$
$=\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}\int_{{}^{d}}u\otimes u\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\mu(\mathrm{d}v)\,\mathrm{d}u$
$=\int_{{}^{d}}\bigl{(}\Psi(v)\otimes\Psi(v)+\mathit{\Sigma}\bigr{)}\,\mu(\mathrm{d}v)\preccurlyeq\kappa_{\Psi}^{2}\mathit{I}_{d}+\mathit{\Sigma},$ (B.3)
which concludes the proof. ∎
It is possible, by using a similar reasoning, to obtain bounds on the moments of $\mathsf{Q}\mathsf{P}\mu$.
Lemma 16.
Let $\mu$ denote a probability measure on ${}^{d}$. Under Assumption H, it holds that
$\bigl{\lvert}\mathcal{M}(\mathsf{Q}\mathsf{P}\mu)\bigr{\rvert}\leqslant\sqrt{\kappa_{\Psi}^{2}+\kappa_{h}^{2}},\qquad$ (B.4)
and
$\min\left\{\frac{\gamma\sigma}{2\kappa_{h}^{2}+\gamma},\frac{\gamma}{2}\right\}\mathit{I}_{d+K}\preccurlyeq\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\preccurlyeq\begin{pmatrix}2\kappa_{\Psi}^{2}\mathit{I}_{d}+2\mathit{\Sigma}&\mathit{0}_{d\times K}\\ \mathit{0}_{K\times d}&2\kappa_{h}^{2}\mathit{I}_{K}+\mathit{\Gamma}\end{pmatrix}.$ (B.5)
Proof.
The inequality (B.4) follows immediately from Assumption H and the fact that
$\mathcal{M}(\mathsf{Q}\mathsf{P}\mu)=\begin{pmatrix}\mathcal{M}(\mathsf{P}\mu)\\ \mathsf{P}\mu[h]\end{pmatrix}.$ (B.8)
For inequality (B.5), let $\phi\colon^{d}\to^{d+K}$ denote the map $\phi(u)=\bigl{(}u,h(u)\bigr{)}$, and let $\phi_{\sharp}$ denote the associated pushforward map on measures. A calculation gives
$\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)=\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)+\begin{pmatrix}\mathit{0}_{d\times d}&\mathit{0}_{d\times K}\\ \mathit{0}_{K\times d}&\mathit{\Gamma}\end{pmatrix}=\begin{pmatrix}\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)&\mathcal{C}^{uy}(\phi_{\sharp}\mathsf{P}\mu)\\ \mathcal{C}^{yu}(\phi_{\sharp}\mathsf{P}\mu)&\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)+\mathit{\Gamma}\end{pmatrix}.$ (B.9)
For any $(a,b)\in^{d}\times^{K}$, it holds that
$2\begin{pmatrix}aa^{\mathsf{T}}&\mathit{0}_{d\times K}\\ \mathit{0}_{K\times d}&bb^{\mathsf{T}}\end{pmatrix}-\begin{pmatrix}a\\ b\end{pmatrix}\otimes\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}a\\ -b\end{pmatrix}\otimes\begin{pmatrix}a\\ -b\end{pmatrix}\succcurlyeq\mathit{0}_{(d+K)\times(d+K)}.$
Therefore, we obtain
$\begin{pmatrix}a\\ b\end{pmatrix}\otimes\begin{pmatrix}a\\ b\end{pmatrix}\preccurlyeq 2\begin{pmatrix}aa^{\mathsf{T}}&\mathit{0}_{d\times K}\\ \mathit{0}_{K\times d}&bb^{\mathsf{T}}\end{pmatrix}$
which enables to deduce, using Lemma 15 and Assumption H, that
$\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d}}\begin{pmatrix}u\\ h(u)\end{pmatrix}\otimes\begin{pmatrix}u\\ h(u)\end{pmatrix}\,\mathsf{P}\mu(u)\,\mathrm{d}u$ (B.14)
$\preccurlyeq 2\int_{{}^{d}}\begin{pmatrix}uu^{\mathsf{T}}&\mathit{0}_{d\times K}\\ \mathit{0}_{K\times d}&h(u)h(u)^{\mathsf{T}}\end{pmatrix}\,\mathsf{P}\mu(u)\,\mathrm{d}u\preccurlyeq 2\begin{pmatrix}\mathcal{\kappa}_{\Psi}^{2}\mathit{I}_{d}+\mathit{\Sigma}&\mathit{0}_{d\times K}\\ \mathit{0}_{K\times d}&\kappa_{h}^{2}\mathit{I}_{K}\end{pmatrix}.$ (B.19)
where we used (B.3) in the last inequality. Combined with (B.9), this inequality leads to the upper bound (B.5). For the lower bound, note that by the Cauchy–Schwarz inequality, it holds for any probability measure $\pi\in\mathcal{P}(^{d}\times^{K})$ and all $(a,b)\in^{d}\times^{K}$
$\left\lvert a^{\mathsf{T}}\mathcal{C}^{uy}(\pi)b\right\rvert$ =×$=\int_{{}^{d}\times^{K}}\left(a^{\mathsf{T}}\bigl{(}u-\mathcal{M}^{u}(\pi)\bigr{)}\right)\left(b^{\mathsf{T}}\bigl{(}y-\mathcal{M}^{y}(\pi)\bigr{)}\right)\pi(\mathrm{d}u\mathrm{d}y)$
$\leqslant\sqrt{a^{\mathsf{T}}\mathcal{C}^{uu}(\pi)a}\,\sqrt{b^{\mathsf{T}}\mathcal{C}^{yy}(\pi)b}.$
Therefore, by Young’s inequality, it holds for all $\varepsilon\in(0,1)$ and for all $(a,b)\in^{d}\times^{K}$, that
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)a-\left(\frac{1}{\varepsilon}-1\right)b^{\mathsf{T}}\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)b$
$\geqslant(1-\varepsilon)a^{\mathsf{T}}\mathit{\Sigma}a-\left(\frac{1}{\varepsilon}-1\right)\kappa_{h}^{2}\lvert b\rvert^{2},$
where we employed (B.1) and the bound $\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)\preccurlyeq\kappa_{H}^{2}\mathit{I}_{K}$ in the last inequality. Using (B.9), we deduce that
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\mathit{\Sigma}a-\left(\frac{1}{\varepsilon}-1\right)\kappa_{h}^{2}\left\lvert b\right\rvert^{2}+b^{\mathsf{T}}\mathit{\Gamma}b$ (B.24)
$\geqslant(1-\varepsilon)\sigma\left\lvert a\right\rvert^{2}+\left(\gamma-\left(\frac{1}{\varepsilon}-1\right)\kappa_{h}^{2}\right)\lvert b\rvert^{2}.$ (B.25)
Letting $\varepsilon$ be such that the coefficient of $\lvert b\rvert^{2}$ is $\gamma/2$, we finally obtain
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}\geqslant\frac{\gamma\sigma}{2\kappa_{h}^{2}+\gamma}\lvert a\rvert^{2}+\frac{\gamma}{2}\lvert b\rvert^{2},$ (B.26)
which concludes the proof. ∎
Remark 17.
A bound sharper than (B.26) can be obtained by letting $\varepsilon$ be such that the coefficients of $\lvert a\rvert^{2}$ and $\lvert b\rvert^{2}$ are equal in (B.25), but this is not necessary for our purposes.
Lemma 18.
For $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ with finite second moments, it holds that
$\bigl{\lvert}\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\bigr{\rvert}$ $\leqslant\frac{1}{2}d_{g}(\mu_{1},\mu_{2}),$
$\bigl{\lVert}\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\bigr{\rVert}$ $\leqslant\left(1+\frac{1}{2}\lvert\mathcal{M}(\mu_{1})+\mathcal{M}(\mu_{2})\rvert\right)\,d_{g}(\mu_{1},\mu_{2}).$
Proof.
Let $m_{i}=\mathcal{M}(\mu_{i})$ and $\mathit{\Sigma}_{i}=\mathcal{C}(\mu_{i})$ for $i=1,2$. Notice that $|2a^{\mathsf{T}}u|\leqslant g(u)$ if $|a|=1$, so
$\lvert m_{1}-m_{2}\rvert=\sup_{\lvert a\rvert=1}\left\lvert a^{\mathsf{T}}\bigl{(}m_{1}-m_{2}\bigr{)}\right\rvert=\sup_{\lvert a\rvert=1}\left\lvert\mu_{1}\bigl{[}a^{\mathsf{T}}u\bigr{]}-\mu_{2}\bigl{[}a^{\mathsf{T}}u\bigr{]}\right\rvert\leqslant\frac{1}{2}d_{g}(\mu_{1},\mu_{2}),$ (B.27)
where the supremum is over the unit sphere in ${}^{n}$, centered at the origin and in the Euclidean distance. Similarly
$\left\lVert\mathit{\Sigma}_{1}-\mathit{\Sigma}_{2}\right\rVert$ $=\sup_{\lvert a\rvert=1}\left\lvert a^{\mathsf{T}}\mathit{\Sigma}_{1}a-a^{\mathsf{T}}\mathit{\Sigma}_{2}a\right\rvert$
$\leqslant\sup_{\lvert a\rvert=1}\left\{\left\lvert\mu_{1}\bigl{[}\lvert a^{\mathsf{T}}u\rvert^{2}\bigr{]}-\mu_{2}\bigl{[}\lvert a^{\mathsf{T}}u\rvert^{2}\bigr{]}\right\rvert+\left\lvert\mu_{1}\bigl{[}a^{\mathsf{T}}u\bigr{]}^{2}-\mu_{2}\bigl{[}a^{\mathsf{T}}u\bigr{]}^{2}\right\rvert\right\}$
$\leqslant d_{g}(\mu_{1},\mu_{2})+\sup_{\lvert a\rvert=1}\left|\mu_{1}\bigl{[}a^{\mathsf{T}}u\bigr{]}+\mu_{2}\bigl{[}a^{\mathsf{T}}u\bigr{]}\right|\left|\mu_{1}\bigl{[}a^{\mathsf{T}}u\bigr{]}-\mu_{2}\bigl{[}a^{\mathsf{T}}u\bigr{]}\right|$
$\leqslant\left(1+\frac{1}{2}\lvert m_{1}+m_{2}\rvert\right)\,d_{g}(\mu_{1},\mu_{2}),$ (B.28)
which concludes the proof. ∎
B.2 Action of $\mathsf{T}_{j}$ on Gaussians
Lemma 19 ($\mathsf{B}_{j}\mathsf{G}=\mathsf{T}_{j}\mathsf{G}$).
Fix $y^{\dagger}_{j+1}\in^{K}$. Let $\pi$ be a Gaussian measure over ×${}^{d}\times^{K}$ with mean and covariance given by
$m=\begin{pmatrix}m_{u}\\ m_{y}\end{pmatrix},\qquad\mathit{S}=\begin{pmatrix}\mathit{S}_{uu}&\mathit{S}_{uy}\\ \mathit{S}_{uy}^{\mathsf{T}}&\mathit{S}_{yy}\end{pmatrix}.$
Recall the pushforward of $\mathscr{T}$ defined in (2.3). Then the probability measure $\mathsf{B}_{j}\pi$ defined in (1.8) coincides with the probability measure $\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi$.
Proof.
For conciseness, we denote $y^{\dagger}=y^{\dagger}_{j+1}$. Using the well known formula for the conditional distribution of a normal random variable, we have that
$\mathsf{B}_{j}\pi=\mathcal{N}\left(m_{u}+S_{uy}S_{yy}^{-1}(y^{\dagger}-m_{y}),S_{uu}-S_{uy}S_{yy}^{-1}S_{uy}^{\mathsf{T}}\right).$ (B.29)
Since $\mathsf{T}_{j}$ is the pushforward under an affine map, it maps Gaussian distributions in $\mathcal{P}(^{d}\times^{K})$ to Gaussian distributions in $\mathcal{P}(^{d})$, and so is sufficient to check that the first and second moments of $\mathsf{B}_{j}\pi$ and $\mathsf{T}_{j}\pi$ coincide. By definition of $\mathsf{T}_{j}$ in (2.3), we have that $U+S_{uy}S_{yy}^{-1}(y^{\dagger}-Y)\sim\mathsf{T}_{j}\pi$ if $(U,Y)\sim\pi$. It follows immediately that the mean under $\mathsf{T}_{j}\pi$ coincides with that under the conditional distribution (B.29). Employing the expression for the mean under $\mathsf{T}_{j}\pi$, we then obtain that the covariance under $\mathsf{T}_{j}\pi$ is given by
$\mathcal{C}(\mathsf{T}_{j}\pi)=\mathbf{E}_{(U,Y)\sim\pi}\biggl{[}\Bigl{(}U-m_{u}+S_{uy}S_{yy}^{-1}\bigl{(}m_{y}-Y\bigr{)}\Bigr{)}\otimes\Bigl{(}U-m_{u}+S_{uy}S_{yy}^{-1}\bigl{(}m_{y}-Y\bigr{)}\Bigr{)}\biggr{]}.$
Developing this expression, we obtain
$\mathcal{C}(\mathsf{T}_{j}\pi)$ $=\mathbf{E}_{(U,Y)\sim\pi}\biggl{[}(U-m_{u})(U-m_{u})^{\mathsf{T}}+(U-m_{u})\bigl{(}m_{y}-Y\bigr{)}^{\mathsf{T}}S_{yy}^{-1}S_{uy}^{\mathsf{T}}$
$\quad\qquad\qquad+S_{uy}S_{yy}^{-1}\bigl{(}m_{y}-Y\bigr{)}(U-m_{u})^{\mathsf{T}}+S_{uy}S_{yy}^{-1}\bigl{(}m_{y}-Y\bigr{)}\bigl{(}m_{y}-Y\bigr{)}^{\mathsf{T}}S_{yy}^{-1}S_{uy}^{\mathsf{T}}\biggr{]}$
$=S_{uu}-S_{uy}S_{yy}^{-1}S_{uy}^{\mathsf{T}}-S_{uy}S_{yy}^{-1}S_{uy}^{\mathsf{T}}+S_{uy}S_{yy}^{-1}S_{uy}^{\mathsf{T}}=S_{uu}-S_{uy}S_{yy}^{-1}S_{uy}^{\mathsf{T}},$
which indeed coincides with the covariance of $\mathsf{B}_{j}\pi$ in (B.29). ∎
Lemma 20.
The maps(3.1) and (3.2) are equivalent.
Proof.
The equivalence follows from Lemma 19 and the operator equality $\mathsf{T}_{j}\mathsf{G}=\mathsf{G}\mathsf{T}_{j}$, with both sides viewed as operators from $\mathcal{P}(^{d}\times^{K})$ to $\mathcal{P}(^{d})$. Since $\mathsf{T}_{j}$ maps Gaussians to Gaussians, the image of both operators is contained in the set $\mathcal{G}(^{d})$ of Gaussian distributions. It is therefore sufficient to check that for any $\pi\in\mathcal{P}(^{d}\times^{K})$, the probability measures $\mathsf{T}_{j}\mathsf{G}\pi$ and $\mathsf{G}\mathsf{T}_{j}\pi$ have the same first and second moments. We saw in the proof of Lemma 19 that the first and second moments of $\mathsf{T}_{j}p$, for any $p\in\mathcal{P}(^{d}\times^{K})$, depend only on the first and second moments of $p$. Therefore the first and second moments of $\mathsf{T}_{j}\pi$ and $\mathsf{T}_{j}\mathsf{G}\pi$ coincide, since the operator $\mathsf{G}$ leaves the first and second moments invariant, and the conclusion follows. ∎
B.3 Stability Results
Lemma 21 (The map $\mathsf{P}$ is globally Lischitz).
Under Assumption H, it holds for all $\mu\in\mathcal{P}(^{d})$ that
$d_{g}(\mathsf{P}\mu,\mathsf{P}\nu)\leqslant\Bigl{(}1+\kappa_{\Psi}^{2}+\operatorname*{tr}(\mathit{\Sigma})\Bigr{)}\,d_{g}(\mu,\nu).$
Proof.
By definition (1.5) of $\mathsf{P}$, it holds that
$\mathsf{P}\mu(u)=\int_{{}^{d}}\frac{\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\,\mu(\mathrm{d}v)=:\int_{{}^{d}}p(v,u)\,\mu(\mathrm{d}v).$
Take any $f\colon^{d}\to$ such that $f\leqslant g$, $g(u)=1+|u|^{2}$. Since $\Psi$ is bounded by assumption,
$b(v):=\int_{{}^{d}}f(u)p(v,u)\,\mathrm{d}u$ $\leqslant\int_{{}^{d}}g(u)p(v,u)\,\mathrm{d}u$
$=1+\left\lvert\Psi(v)\right\rvert^{2}+\operatorname*{tr}(\mathit{\Sigma})\leqslant 1+\kappa_{\Psi}^{2}+\operatorname*{tr}(\mathit{\Sigma}).$
Therefore, using Fubini’s theorem, we have
$\Bigl{\lvert}\mathsf{P}\mu[f]-\mathsf{P}\nu[f]\Bigr{\rvert}$ $=\biggl{\lvert}\int_{{}^{d}}\left(\int_{{}^{d}}f(u)p(v,u)\,\mathrm{d}u\right)\bigl{(}\mu(\mathrm{d}v)-\nu(\mathrm{d}v)\bigr{)}\biggr{\rvert}$
$=\Bigl{\lvert}\mu[b]-\nu[b]\Bigr{\rvert}\leqslant\Bigl{(}1+\kappa_{\Psi}^{2}+\operatorname*{tr}(\mathit{\Sigma})\Bigr{)}d_{g}(\mu,\nu),$
which concludes the proof. ∎
Lemma 22 (The map $\mathsf{Q}$ is globally Lipschitz).
Under Assumption H, it holds for any $\mu,\nu\in\mathcal{P}(^{d})$ that
$d_{g}(\mathsf{Q}\mu,\mathsf{Q}\nu)\leqslant\Bigl{(}1+\kappa_{h}^{2}+\operatorname*{tr}(\mathit{\Gamma})\Bigr{)}d_{g}(\mu,\nu).$ (B.30)
Proof.
Let us take $f\colon^{d}\times^{K}\to$ such that $f\leqslant g$ and introduce the operator $\Pi$ given by
$\Pi f(u)=\int_{{}^{K}}f(u,y)\,\mathcal{N}\bigl{(}h(u),\mathit{\Gamma}\bigr{)}(\mathrm{d}y).$
Clearly $\Pi f(u)\leqslant\Pi g(u)$, and it holds that
$\Pi g(u)$ $=\int_{{}^{K}}\left(1+\lvert u\rvert^{2}+\lvert y\rvert^{2}\right)\,\mathcal{N}\bigl{(}h(u),\mathit{\Gamma}\bigr{)}(\mathrm{d}y)=1+\lvert u\rvert^{2}+\int_{{}^{K}}\lvert y\rvert^{2}\mathcal{N}\bigl{(}h(u),\mathit{\Gamma}\bigr{)}(\mathrm{d}y)$
$=1+\lvert u\rvert^{2}+\bigl{\lvert}h(u)\bigr{\rvert}^{2}+\operatorname*{tr}(\mathit{\Gamma})\leqslant\Bigl{(}1+\kappa_{h}^{2}+\operatorname*{tr}(\mathit{\Gamma})\Bigr{)}\left(1+\lvert u\rvert^{2}\right).$
Therefore
$\Bigl{\lvert}\mathsf{Q}\mu[f]-\mathsf{Q}\nu[f]\Bigr{\rvert}=\Bigl{\lvert}\mu[\Pi f]-\nu[\Pi f]\Bigr{\rvert}\leqslant\Bigl{(}1+\kappa_{h}^{2}+\operatorname*{tr}(\mathit{\Gamma})\Bigr{)}d_{g}(\mu,\nu),$
and we obtain (B.30). ∎
Lemma 23.
Under Assumption H, there exists $C_{\mathsf{B}}=C_{\mathsf{B}}(\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$ such that for any probability measure $\mu\in\mathcal{P}(^{d})$, it holds that
$\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)\leqslant C_{\mathsf{B}}d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
Proof.
For conciseness, we denote $y^{\dagger}=y^{\dagger}_{j+1}$. Let us introduce the $y$-marginal densities
$\alpha_{\mu}(y):=\int_{{}^{d}}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\,,\qquad\beta_{\mu}(y):=\int_{{}^{d}}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\,.$
Then
$d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)=\int_{{}^{d}}\left(1+|u|^{2}\right)\left|\frac{\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})}{\alpha_{\mu}(y^{\dagger})}-\frac{\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})}{\beta_{\mu}(y^{\dagger})}\right|\,\mathrm{d}u$
$\qquad\qquad\leqslant\frac{1}{\alpha_{\mu}(y^{\dagger})}\int_{{}^{d}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})-\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})\bigr{\rvert}\,\mathrm{d}u$
$\qquad\qquad\qquad+\left|\frac{\alpha_{\mu}(y^{\dagger})-\beta_{\mu}(y^{\dagger})}{\alpha_{\mu}(y^{\dagger})\beta_{\mu}(y^{\dagger})}\right|\,\int_{{}^{d}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})\,\mathrm{d}u.$ (B.31)
Step 1: bounding $\alpha_{\mu}(y^{\dagger})$ and $\beta_{\mu}(y^{\dagger})$ from below
The marginal distribution $\alpha_{\mu}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ is Gaussian with covariance matrix
[…]
which is bounded from below by $\mathit{\Gamma}$ and from above by $\mathit{\Gamma}+\kappa_{h}^{2}\mathit{I}_{K}$, in view of Assumption H. (We again use the fact that $aa^{\mathsf{T}}\preccurlyeq(a^{\mathsf{T}}a)\mathit{I}_{d}$ for any vector $a\in^{d}$, by the Cauchy-Schwarz inequality.) Assumption H also implies that the mean of $\alpha_{\mu}$ is bounded from above in norm by $\kappa_{h}$. Therefore, it holds that
$\forall y\in^{K},\qquad\alpha_{\mu}(y)\geqslant\frac{\exp\Bigl{(}-\frac{1}{2}(|y|+\kappa_{h})^{2}\left\lVert\mathit{\Gamma}^{-1}\right\rVert\Bigr{)}}{\sqrt{(2\pi)^{K}\det(\mathit{\Gamma}+\kappa_{h}^{2}\mathit{I}_{K})}}.$ (B.33)
The function $\beta_{\mu}$ can be bounded from below independently of $\mu$ in a similar manner. Indeed, it holds under Assumption H that for all $y\in^{K}$,
$\beta_{\mu}(y)$ $=\int_{{}^{d}}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u=\int_{{}^{d}}\frac{\exp\left(-\frac{1}{2}\bigl{(}y-h(u)\bigr{)}^{\mathsf{T}}\mathit{\Gamma}^{-1}\bigl{(}y-h(u)\bigr{)}\right)}{\sqrt{(2\pi)^{K}\det(\mathit{\Gamma})}}\mathsf{P}\mu(u)\,\mathrm{d}u$
$\geqslant\frac{\exp\Bigl{(}-\frac{1}{2}(|y|+\kappa_{h})^{2}\left\lVert\mathit{\Gamma}^{-1}\right\rVert\Bigr{)}}{\sqrt{(2\pi)^{K}\det(\mathit{\Gamma})}}.$ (B.34)
Step 2: bounding the first term in (B.31)
For fixed $u\in^{d}$, the functions $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)$ and $y\mapsto\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)$ are Gaussians up to constant factors. The covariance matrix of the former is $\mathit{\Gamma}$ and, using the formula for the covariance of the conditional distribution of a Gaussian, we calculate that the covariance of the latter is given by
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)-\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ (B.35)
Since $\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ is positive semi-definite, it follows from Lemma 16 that the matrix (B.35) is bounded from above by $2\kappa_{h}^{2}\mathit{I}_{K}+\mathit{\Gamma}$. Then, using the same notation as in (B.9), we obtain that
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)-\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)\\ =\mathit{\Gamma}+\Bigl{(}\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)-\mathcal{C}^{yu}(\phi_{\sharp}\mathsf{P}\mu)\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\phi_{\sharp}\mathsf{P}\mu)\Bigr{)}\succcurlyeq\mathit{\Gamma},$ (B.36)
where the inequality holds because, being the Schur complement of the block $\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)$ of the matrix $\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$, the bracketed term is positive semi-definite. Therefore, the matrix (B.35) is bounded from below by $\mathit{\Gamma}$, and so the integral in the first term of (B.31) may be bounded from above by using Lemma 12 in Appendix A with parameter $\alpha=\alpha(\kappa_{h},\mathit{\Gamma})$, which gives
$\int_{{}^{d}}\left(1+|u|^{2}\right)\left|\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\right|\mathrm{d}u$
$\qquad\leqslant C\int_{{}^{K}}\int_{{}^{d}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
$\qquad\leqslant C\int_{{}^{K}}\int_{{}^{d}}\left(1+|u|^{2}+|y|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
$\qquad=Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$ (B.37)
for an appropriate constant $C$ depending only on $\mathit{\Gamma}$ and $\kappa_{h}$.
Step 3: bounding the second term in (B.31)
By (B.37), we have
$\bigl{\lvert}\alpha_{\mu}(y)-\beta_{\mu}(y)\bigr{\rvert}$ $=\left|\int_{{}^{d}}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\right|$
$\leqslant\int_{{}^{d}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
On the other hand, since $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)/\mathsf{P}\mu(u)$ is a Gaussian density with covariance $\mathit{\Gamma}$, which is bounded uniformly from above by $\bigl{(}(2\pi)^{K}\det(\mathit{\Gamma})\bigr{)}^{-1/2}$, it holds that
$\int_{{}^{d}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$ $=\beta_{\mu}(y)+\int_{{}^{d}}|u|^{2}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$
$\leqslant\beta_{\mu}(y)+\int_{{}^{d}}\frac{|u|^{2}\mathsf{P}\mu(u)}{\sqrt{(2\pi)^{K}\det(\Gamma})}\,\mathrm{d}u$
$=\beta_{\mu}(y)+\frac{\operatorname*{tr}\bigl{(}\mathcal{C}(\mathsf{P}\mu)\bigr{)}+\left\lvert\mathcal{M}(\mathsf{P}\mu)\right\rvert^{2}}{\sqrt{(2\pi)^{K}\det(\mathit{\Gamma}})}.$
By Lemma 15, the first and second moments of $\mathsf{P}\mu$ are bounded from above by a constant depending only on $\kappa_{\Psi}$ and $\mathit{\Sigma}$.
Step 4: concluding the proof
Combining the above inequalities, we deduce that
$d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)\leqslant\frac{C(\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})}{\alpha_{\mu}(y^{\dagger}_{j+1})}\left(1+\frac{1}{\beta_{\mu}(y^{\dagger}_{j+1})}\right)d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
Using (B.33) and (B.34) together with the uniform bound $\kappa_{y}$ on the data (Assumption H) gives the conclusion. ∎
To conclude this section, we show a stability result for $\mathsf{T}_{j}$.
Lemma 24 (Stability result for the mean field map $\mathsf{T}_{j}$).
Suppose that Assumption H is satisfied and that $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$. Then, for all $R\geqslant 1$, there is $L_{\mathsf{T}}=L_{\mathsf{T}}(R,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\mathit{\Sigma},\mathit{\Gamma})$ such that for all $\pi\in\mathcal{P}_{R}(^{d}\times^{K})$ and $\mu\in\mathcal{P}(^{d})$, it holds that
$\forall j\in,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p),\qquad p:=\mathsf{Q}\mathsf{P}\mu.$
Proof.
By Lemma 16, the probability measure $\mathsf{Q}\mathsf{P}\mu$ belongs to $\mathcal{P}_{\widetilde{R}}(^{d}\times^{K})$ for some $\widetilde{R}\geqslant 1$. Let us introduce
$r=\max\Bigl{\{}R,\widetilde{R},\kappa_{y}\Bigr{\}}$
and denote by $\mathscr{T}^{\pi}$ and $\mathscr{T}^{p}$ the affine maps corresponding to evaluation of covariance information at $\pi$ and $p=\mathsf{Q}\mathsf{P}\mu$. Specifically,
$\mathscr{T}^{\pi}(u,y)$ $=u+\mathit{A}_{\pi}(y^{\dagger}-y),\qquad$ $\mathit{A}_{\pi}:=\mathit{K}_{uy}\mathit{K}_{yy}^{-1},$
$\mathscr{T}^{p}(u,y)$ $=u+\mathit{A}_{p}(y^{\dagger}-y),\qquad$ $\mathit{A}_{p}:=\mathit{S}_{uy}\mathit{S}_{yy}^{-1},$
where $\mathit{K}=\mathcal{C}(\pi)$, $\mathit{S}=\mathcal{C}(p)$ and $y^{\dagger}=y^{\dagger}_{j+1}$. By the triangle inequality, we have
$d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant d_{g}(\mathscr{T}^{\pi}_{\sharp}\pi,\mathscr{T}^{\pi}_{\sharp}p)+d_{g}(\mathscr{T}^{\pi}_{\sharp}p,\mathscr{T}^{p}_{\sharp}p).$ (B.38)
Before separately bounding each term on the right-hand side, we obtain simple inequalities that will be helpful later in the proof. It holds that
$\lVert\mathit{A}_{\pi}\rVert\leqslant\lVert\mathit{K}_{uy}\rVert\lVert\mathit{K}_{yy}^{-1}\rVert\leqslant\lVert\mathit{K}\rVert\lVert\mathit{K}^{-1}\rVert\leqslant r^{4},$ (B.39)
and similarly for $\mathit{A}_{p}$. Here we used that the matrix norm (induced by the Euclidean vector norm) of any submatrix is bounded from above by the norm of the original matrix. Using this bound again, we obtain, assuming $r>1$ without any loss of generality,
$\lVert\mathit{A}_{\pi}-\mathit{A}_{p}\rVert=\lVert\mathit{K}_{uy}\mathit{K}_{yy}^{-1}-\mathit{S}_{uy}\mathit{S}_{yy}^{-1}\rVert$ $\leqslant\lVert(\mathit{K}_{uy}-\mathit{S}_{uy})\mathit{K}_{yy}^{-1}\rVert+\lVert\mathit{S}_{uy}\left(\mathit{K}_{yy}^{-1}-\mathit{S}_{yy}^{-1}\right)\rVert$
$\leqslant r^{2}\Bigl{(}\lVert\mathit{K}_{uy}-\mathit{S}_{uy}\rVert+\lVert\mathit{K}_{yy}^{-1}-\mathit{S}_{yy}^{-1}\rVert\Bigr{)}$
$=r^{2}\Bigl{(}\lVert\mathit{K}_{uy}-\mathit{S}_{uy}\rVert+\lVert\mathit{K}_{yy}^{-1}(\mathit{S}_{yy}-\mathit{K}_{yy})\mathit{S}_{yy}^{-1}\rVert\Bigr{)}$
$\leqslant r^{2}\Bigl{(}\lVert\mathit{K}_{uy}-\mathit{S}_{uy}\rVert+\lVert\mathit{K}_{yy}^{-1}\rVert\lVert\mathit{S}_{yy}-\mathit{K}_{yy}\rVert\lVert\mathit{S}_{yy}^{-1}\rVert\Bigr{)}$
$\leqslant 2r^{6}\lVert\mathit{K}-\mathit{S}\rVert\leqslant 2r^{6}(1+2r)\,d_{g}(\pi,p),$ (B.40)
where we used in the last line the inequality (B.28) from Lemma 18.
Bounding the first term in (B.38)
Let $f\colon^{d}\to$ denote a function satisfying $\lvert f\rvert\leqslant g$. It follows from the definition of the pushforward measure that
$\bigl{\lvert}\mathscr{T}^{\pi}_{\sharp}\pi[f]-\mathscr{T}^{\pi}_{\sharp}p[f]\bigr{\rvert}=\bigl{\lvert}\pi[f\circ\mathscr{T}^{\pi}]-p[f\circ\mathscr{T}^{\pi}]\bigr{\rvert}.$
For all $(u,y)\in^{d}\times^{K}$, it holds that
$\bigl{\lvert}f\circ\mathscr{T}^{\pi}(u,y)\bigr{\rvert}$ $=\left\lvert f\Bigl{(}u+\mathit{A}_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}\right\rvert\leqslant g\Bigl{(}u+\mathit{A}_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}$
$=1+\bigl{\lvert}u+\mathit{A}_{\pi}(y^{\dagger}-y)\bigr{\rvert}^{2}\leqslant 1+3\lvert u\rvert^{2}+3\lvert\mathit{A}_{\pi}y^{\dagger}\rvert^{2}+3\lvert\mathit{A}_{\pi}y\rvert^{2}$
$\leqslant 3\left(1+\lvert\mathit{A}_{\pi}y^{\dagger}\rvert^{2}\right)\max\left\{1,\lVert A_{\pi}\rVert^{2}\right\}g(u,y).$
Therefore, using (B.39) we deduce that
$d_{g}(\mathscr{T}^{\pi}_{\sharp}\pi,\mathscr{T}^{\pi}_{\sharp}p)\leqslant 3\left(1+r^{10}\right)r^{8}\,d_{g}(\pi,p).$ (B.41)
Bounding the second term in (B.38)
Let $f\colon^{d}\to$ again denote a function satisfying the inequality $\lvert f\rvert\leqslant g$, with $g(u)=1+\lvert u\rvert^{2}$, It holds that
$\Bigl{\lvert}\mathscr{T}^{\pi}_{\sharp}p[f]-\mathscr{T}^{p}_{\sharp}p[f]\Bigr{\rvert}=\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}]-p[f\circ\mathscr{T}^{p}]\Bigr{\rvert}=\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}.$
The right-hand side may be rewritten more explicitly as
$\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}$ $=\left\lvert\int_{{}^{K}}\int_{{}^{d}}f\left(u+\mathit{A}_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\right)-f\left(u+\mathit{A}_{p}\bigl{(}y^{\dagger}-y\bigr{)}\right)\,p(u,y)\,\mathrm{d}u\,\mathrm{d}y\right\rvert.$
We apply a change of variable in order rewrite the integral in this identity as
$\int_{{}^{K}}\int_{{}^{d}}\bigl{(}f(u+\mathit{A}_{\pi}z)-f(u+\mathit{A}_{p}z)\bigr{)}\,p(u,y^{\dagger}-z)\,\mathrm{d}u\,\mathrm{d}z$
$=\int_{{}^{K}}\int_{{}^{d}}f(v)\Bigl{(}p(v-\mathit{A}_{\pi}z,y^{\dagger}-z)-p(v-\mathit{A}_{p}z,y^{\dagger}-z)\Bigr{)}\mathrm{d}v\,\mathrm{d}z.$ (B.42)
It follows from the technical Lemma 14 proved in Appendix A that
$\Bigl{\lvert}p(v-\mathit{A}_{\pi}z,y^{\dagger}-z)-p(v-\mathit{A}_{p}z,y^{\dagger}-z)\Bigr{\rvert}$
$\qquad\leqslant C\lvert\mathit{A}_{\pi}z-\mathit{A}_{p}z\rvert\exp\left(-\frac{1}{4}\left(\left\lvert y^{\dagger}-z\right\rvert_{\mathit{\Gamma}}^{2}+\min\left\{\left\lvert v-\mathit{A}_{\pi}z\right\rvert_{K}^{2},\left\lvert v-\mathit{A}_{p}z\right\rvert_{\mathit{K}}^{2}\right\}\right)\right).$
Using this inequality, we first bound the inner integral in (B.42) for fixed $z\in^{K}$. Keeping only the terms that depend on $v$, we obtain that
$\int_{{}^{d}}\lvert f(v)\rvert\exp\left(-\frac{1}{4}\min\left\{\left\lvert v-\mathit{A}_{\pi}z\right\rvert_{\mathit{K}}^{2},\left\lvert v-\mathit{A}_{p}z\right\rvert_{\mathit{K}}^{2}\right\}\right)\,\mathrm{d}v$
$\qquad=\int_{{}^{d}}\lvert f(v)\rvert\max\left\{\exp\left(-\frac{1}{4}\left\lvert v-\mathit{A}_{\pi}z\right\rvert_{\mathit{K}}^{2}\right),\exp\left(-\frac{1}{4}\left\lvert v-\mathit{A}_{p}z\right\rvert_{\mathit{K}}^{2}\right)\right\}\,\mathrm{d}v$
$\qquad\leqslant\int_{{}^{d}}\lvert f(v)\rvert\exp\left(-\frac{1}{4}\left\lvert v-\mathit{A}_{\pi}z\right\rvert_{\mathit{K}}^{2}\right)\mathrm{d}v+\int_{{}^{d}}\lvert f(v)\rvert\exp\left(-\frac{1}{4}\left\lvert v-\mathit{A}_{p}z\right\rvert_{\mathit{K}}^{2}\right)\mathrm{d}v.$
Since $\lvert f(v)\rvert\leqslant 1+\lvert v\rvert^{2}$, it is clear that this expression is bounded from above by
$C\Bigl{(}1+\lvert\mathit{A}_{\pi}z\rvert^{2}+\lvert\mathit{A}_{p}z\rvert^{2}\Bigr{)}\leqslant Cr^{8}\left(1+\lvert z\rvert^{2}\right),$
where we used (B.39) in the last inequality. The remaining integral in the $z$ direction can be bounded similarly:
$d_{g}(\mathscr{T}^{\pi}_{\sharp}p,\mathscr{T}^{p}_{\sharp}p)$ $\leqslant Cr^{8}\int_{{}^{K}}\left(1+\lvert z\rvert^{2}\right)\lvert\mathit{A}_{\pi}z-\mathit{A}_{p}z\rvert\exp\left(-\frac{1}{4}\left(\left\lvert y^{\dagger}-z\right\rvert_{\mathit{\Gamma}}^{2}\right)\right)\,\mathrm{d}z$
$\leqslant Cr^{11}\lVert\mathit{A}_{\pi}-\mathit{A}_{p}\rVert\leqslant Cr^{18}d_{g}(\pi,p),$ (B.43)
where we used the bound (B.40) in the last inequality. Combining (B.41) and (B.43), we deduce the statement. ∎
B.4 Additional Stability Results for the Gaussian Projected Filter
Next, we show the local Lipschitz continuity of the Gaussian projection map $\mathsf{G}$. To this end, we begin by proving the following result. (The proof uses an approach similar to that employed in [16] for obtaining an upper bound on the distance between Gaussians in the usual total variation metric.)
Lemma 25.
Let $\mu_{1}=\mathcal{N}(m_{1},\mathit{\Sigma}_{1})$ and $\mu_{2}=\mathcal{N}(m_{2},\mathit{\Sigma}_{2})$ with $\mathit{\Sigma}_{1},\mathit{\Sigma}_{2}\in^{n\times n}$ symmetric and positive definite. It holds that
$d_{g}\bigl{(}\mu_{1},\mu_{2}\bigr{)}\leqslant\sqrt{\mu_{1}\left[g^{2}\right]+\mu_{2}\left[g^{2}\right]}\left(3\left\lVert\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n}\right\rVert_{\rm F}+\lvert m_{1}-m_{2}\rvert_{\mathit{\Sigma}_{2}}\right),$ (B.44)
where $\lVert\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\rVert_{\rm F}$ is the Frobenius matrix norm.
Proof.
We denote by $(\lambda_{i})_{1\leqslant i\leqslant n}$ the eigenvalues of $\mathit{\Sigma}_{1}^{-1}\mathit{\Sigma}_{2}$. Being the product of two symmetric positive definite matrices, $\mathit{\Sigma}_{1}^{-1}\mathit{\Sigma}_{2}$ is real diagonalizable with real and positive eigenvalues. We note that
$d_{g}(\mu_{1},\mu_{2})\leqslant\mu_{1}[g]+\mu_{2}[g]\leqslant\sqrt{\mu_{1}[g^{2}]}+\sqrt{\mu_{2}[g^{2}]}\leqslant\sqrt{2}\sqrt{\mu_{1}[g^{2}]+\mu_{2}[g^{2}]}.$
Now assume $\max\{\lambda_{j}:1\leqslant j\leqslant n\}\geqslant 2$. In this case there is integer $k$ such that $\lambda_{k}\geqslant 2$ and so
$3\lVert\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n}\rVert_{\rm F}=3\sqrt{\sum_{i=1}^{n}\lvert\lambda_{i}^{-1}-1\rvert^{2}}\geqslant 3\left\lvert\lambda_{k}^{-1}-1\right\rvert\geqslant\frac{3}{2}\geqslant\sqrt{2}.$
Combining the previous two inequalities proves (B.44) in the case where
$\max\{\lambda_{j}:1\leqslant j\leqslant n\}\geqslant 2.$
We assume from now on that $20<\lambda_{i}<$ 2 for all $i\in$. Employing the same reasoning as in [2, Lemma 3.1], we can prove
$d_{g}(\mu_{1},\mu_{2})^{2}\leqslant 2\Bigl{(}\mu_{1}\bigl{[}g^{2}\bigr{]}+\mu_{2}\bigl{[}g^{2}\bigr{]}\Bigr{)}\,\mathrm{KL}(\mu_{2}\|\mu_{1}),$ (B.45)
where $\mathrm{KL}(\mu_{2}\|\mu_{1})$ is the Kullback–Leibler (KL) divergence of $\mu_{2}$ from $\mu_{1}$, given by
$\mathrm{KL}(\mu_{1}\|\mu_{2}):=\int_{{}^{d}}\frac{\mathrm{d}\mu_{1}}{\mathrm{d}\mu_{2}}\,\log\left(\frac{\mathrm{d}\mu_{1}}{\mathrm{d}\mu_{2}}\right)\,\mathrm{d}\mu_{2}.$ (B.46)
The proof of the inequality (B.45) is presented in Lemma 11 in Appendix A for completeness. The KL divergence between two Gaussians has a closed expression, which we rewrite in terms of the function $f_{1}(x):=x^{-1}-1+\log x$,
$\mathrm{KL}(\mu_{2}\|\mu_{1})$ $=\frac{1}{2}\left(\operatorname*{tr}(\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n})+(m_{1}-m_{2})^{\mathsf{T}}\mathit{\Sigma}_{2}^{-1}(m_{1}-m_{2})-\log\det(\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1})\right)$
$=\frac{1}{2}\left(\sum_{i=1}^{n}(\lambda_{i}^{-1}-1)+(m_{1}-m_{2})^{\mathsf{T}}\mathit{\Sigma}_{2}^{-1}(m_{1}-m_{2})+\sum_{i=1}^{n}\log\lambda_{i}\right)$
$=\frac{1}{2}\left(\sum_{i=1}^{n}f_{1}(\lambda_{i})+\lvert m_{1}-m_{2}\rvert_{\mathit{\Sigma}_{2}}^{2}\right).$
The function $f_{1}$ is pointwise bounded from above by $f_{2}(x):=(1-x^{-1})^{2}$ for $x\in(0,2)$. To see this consider the function ${\varphi}(x)=f_{2}(x)-f_{1}(x)$. It suffices to note that $\varphi(1)=0$ and
$\varphi^{\prime}(x)=-\frac{1}{x}\left(1-\frac{1}{x}\right)\left(1-\frac{2}{x}\right).$
Since $\varphi^{\prime}(x)<0$ for $x\in(0,1)$ and $\varphi^{\prime}(x)>0$ for $x\in(1,2)$, the desired upper bound on $f_{1}(x)$ by $f_{2}(x)$ for $x\in(0,2)$ follows. Therefore, since
$\sum_{i=1}^{n}f_{1}(\lambda_{i})\leqslant\sum_{i=1}^{n}f_{2}(\lambda_{i})=\lVert\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n}\rVert_{\rm F}^{2},$
we have
$\mathrm{KL}(\mu_{2}\|\mu_{1})$ $\leqslant\frac{1}{2}\Bigl{(}\lVert\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n}\rVert_{\rm F}^{2}+\lvert m_{1}-m_{2}\rvert_{\mathit{\Sigma}_{2}}^{2}\Bigr{)}$
$\leqslant\frac{1}{2}\Bigl{(}\lVert\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n}\rVert_{\rm F}+\lvert m_{1}-m_{2}\rvert_{\mathit{\Sigma}_{2}}\Bigr{)}^{2}.$
Consequently, we deduce from (B.45) that
$d_{g}(\mu_{1},\mu_{2})$ $\leqslant\sqrt{\mu_{1}\bigl{[}g^{2}\bigr{]}+\mu_{2}\bigl{[}g^{2}\bigr{]}}\Bigl{(}\lVert\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n}\rVert_{\rm F}+\lvert m_{1}-m_{2}\rvert_{\mathit{\Sigma}_{2}}\Bigr{)},$
which concludes the proof. ∎
Lemma 26.
For all $R\geqslant 1$, there is $L_{\mathsf{G}}=L_{\mathsf{G}}(R,n)$ such that for all $\mu_{1},\mu_{2}\in\mathcal{P}_{R}(^{n})$, the following inequality holds:
$d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})\leqslant L_{\mathsf{G}}(R,n)\,d_{g}(\mu_{1},\mu_{2}),$
Before proving this lemma we make two remarks on the form of the result.
Remark 27.
Lemmas 21 and 22 on the Lipschitz continuity of $\mathsf{P}$ and $\mathsf{Q}$ do not depend on the specific choice of the function $g$ in Definition 1; they hold true also for the usual total variation distance. In contrast, in the proof of Lemma 26 the specific choice of $g$ is used in order to control differences between moments by means of $d_{g}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$.
Remark 28.
One may wonder whether the non-unit Lipschitz constant in Lemma 26 is inherent or just an artifact of the method of proof. A simple example shows that the Gaussian projection operator $\mathsf{G}$ is not globally Lipschitz continuous. Specifically, consider the probability distributions $\mu_{1}=\mathcal{N}(0,\varepsilon)$ with a variance $\varepsilon\leqslant 1/2$ and
$\mu_{2}=\varepsilon\delta_{-1}+(1-2\varepsilon)\mu_{1}+\varepsilon\delta_{1}.$
By definition of $d_{g}$, it follows that
$d_{g}(\mu_{1},\mu_{2})$ $\leqslant\varepsilon g(-1)+2\varepsilon\mu_{1}[g]+\varepsilon g(1)$
$=2\varepsilon+2\varepsilon\left(1+\varepsilon\right)+2\varepsilon\leqslant 8\varepsilon.$ (B.47)
On the other hand, $\mathsf{G}\mu_{2}=\mathcal{N}\bigl{(}0,2\varepsilon+(1-2\varepsilon)\varepsilon\bigr{)}$, and the variance of this Gaussian is bounded from below by $3\varepsilon-2\varepsilon^{2}\geqslant 2\varepsilon$. Consequently, it holds that
$d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})\geqslant d_{1}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})\geqslant d_{1}\bigl{(}\mathcal{N}(0,\varepsilon),\mathcal{N}(0,2\varepsilon)\bigr{)}=d_{1}\bigl{(}\mathcal{N}(0,1),\mathcal{N}(0,2)\bigr{)},$
where $d_{1}$ is the usual total variation metric. The right-hand side of this inequality does not depend on $\varepsilon$, and so
$\frac{d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})}{d_{g}(\mu_{2},\mu_{2})}\xrightarrow[\varepsilon\to 0]{}\infty.$
This proves that $\mathsf{G}$ is not globally Lipschitz.
Proof of Lemma 26.
Let $m_{i}=\mathcal{M}(\mu_{i})$ and $\mathit{\Sigma}_{i}=\mathcal{C}(\mu_{i})$ for $i=1,2$. By Lemma 18, it holds that
$\lvert m_{1}-m_{2}\rvert_{\mathit{\Sigma}_{2}}\leqslant\frac{1}{2}\sqrt{\lVert\mathit{\Sigma}_{2}^{-1}\rVert}\,d_{g}(\mu_{1},\mu_{2}).$
In addition, also using Lemma 18, and the facts that $\lVert A\rVert_{\rm F}\leqslant\sqrt{n}\,\lVert A\rVert$ and $\lVert AB\rVert\leqslant\lVert A\rVert\lVert B\rVert$ for any matrices $A,B\in^{n\times n}$,
$\lVert\mathit{\Sigma}_{2}^{-1}\mathit{\Sigma}_{1}-I_{n}\rVert_{\rm F}\leqslant\sqrt{n}\lVert\mathit{\Sigma}_{2}^{-1}\rVert\left(1+\frac{1}{2}\lvert m_{1}+m_{2}\rvert\right)\,d_{g}(\mu_{1},\mu_{2}).$
Lemma 25 then gives
$d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})$ $\leqslant\sqrt{\mathsf{G}\mu_{1}\left[g^{2}\right]+\mathsf{G}\mu_{2}\left[g^{2}\right]}\left(3\sqrt{n}\left(1+\frac{1}{2}\lvert m_{1}+m_{2}\rvert\right)\left\lVert\mathit{\Sigma}_{2}^{-1}\right\rVert+\frac{1}{2}\sqrt{\left\lVert\mathit{\Sigma}_{2}^{-1}\right\rVert}\right)d_{g}(\mu_{1},\mu_{2})$
$\leqslant\sqrt{\mathsf{G}\mu_{1}[g^{2}]+\mathsf{G}\mu_{2}[g^{2}]}\left(3\sqrt{n}\left(1+\frac{1}{2}\lvert m_{1}+m_{2}\rvert\right)+1\right)\left(1+\left\lVert\mathit{\Sigma}_{2}^{-1}\right\rVert\right)d_{g}(\mu_{1},\mu_{2}).$
Clearly $\mathsf{G}\mu_{i}\left[g^{2}\right]\leqslant C\bigl{(}1+\lvert m_{i}\rvert^{4}+\lVert\mathit{\Sigma}_{i}\rVert^{2}\bigr{)}$ for our choice of $g$, for an appropriate constant $C$, and so we conclude by use of the Young inequality that
$d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})\leqslant C\left(1+\lvert m_{1}\rvert^{3}+\lvert m_{2}\rvert^{3}+\left\lVert\mathit{\Sigma}_{1}\right\rVert^{3/2}+\left\lVert\mathit{\Sigma}_{2}\right\rVert^{3/2}\right)\bigl{(}1+\left\lVert\mathit{\Sigma}_{2}^{-1}\right\rVert\bigr{)}\,d_{g}(\mu_{1},\mu_{2}),$
for another constant $C$ depending only on $n$; this completes the proof. ∎
We end the subsection by showing local Lipschitz continuity of the conditioning operator $\mathsf{B}_{j}$ over the set $\mathcal{G}(^{d}\times^{K})$.
Lemma 29.
For all $R\geqslant 1$, there exists a constant $L_{\mathsf{B}}=L_{\mathsf{B}}(R,\kappa_{y})$ such that for all Gaussian probability measures $\pi,p\in\mathcal{G}_{R}(^{d}\times^{K})$,
$\forall j\in,\qquad d_{g}\bigl{(}\mathsf{B}_{j}\pi,\mathsf{B}_{j}p\bigr{)}\leqslant L_{\mathsf{B}}(R,\kappa_{y})\,d_{g}\bigl{(}\pi,p\bigr{)}.$ (B.48)
Proof.
Let $\pi=\mathcal{N}({\tau},\mathit{\Upsilon})$ and $p=\mathcal{N}({t},\mathit{U})$. We use the shorthand notation $y^{\dagger}=y^{\dagger}_{j+1}$. It is well known that if $Z\sim\mathcal{N}({\tau},\mathit{\Upsilon})$, for a vector $\tau\in^{d+K}$ and a matrix $\mathit{\Upsilon}\in^{(d+K)\times(d+K)}$, then the conditional distribution of $Z_{u}$ given that $Z_{y}=y^{\dagger}$ is given by
$Z_{u}|Z_{y}=y^{\dagger}\sim\mathcal{N}\bigl{(}{\tau}_{u}+\mathit{\Upsilon}_{uy}\mathit{\Upsilon}_{yy}^{-1}(y^{\dagger}-{\tau}_{y}),\mathit{\Upsilon}_{uu}-\mathit{\Upsilon}_{uy}\mathit{\Upsilon}_{yy}^{-1}\mathit{\Upsilon}_{yu}\bigr{)}.$ (B.49)
Denoting $\mathsf{B}_{j}\pi=\mathcal{N}(\lambda,\mathit{\Delta})$ and $\mathsf{B}_{j}p=\mathcal{N}(\ell,\mathit{D})$, we have
$\lambda-\ell$ $={\tau}_{u}-{t}_{u}+\mathit{\Upsilon}_{uy}\mathit{\Upsilon}_{yy}^{-1}(y^{\dagger}-{\tau}_{y})-\mathit{U}_{uy}\mathit{U}_{yy}^{-1}(y^{\dagger}-{t}_{y})$
$={\tau}_{u}-{t}_{u}+(\mathit{\Upsilon}_{uy}-\mathit{U}_{uy})\mathit{\Upsilon}_{yy}^{-1}(y^{\dagger}-{\tau}_{y})$
$\qquad+\mathit{U}_{uy}\mathit{U}_{yy}^{-1}(\mathit{U}_{yy}-\mathit{\Upsilon}_{yy})\mathit{\Upsilon}_{yy}^{-1}(y^{\dagger}-{\tau}_{y})-\mathit{U}_{uy}\mathit{U}_{yy}^{-1}(\tau_{y}-t_{y}).$
Since the 2-norm of any submatrix is bounded from above by the 2-norm of the matrix that contains it, we deduce
$\bigl{\lvert}\lambda-\ell\bigr{\rvert}\leqslant 2R^{4}\left\lvert{\tau}-{t}\right\rvert+2R^{6}(R+\kappa_{y})\lVert\mathit{\Upsilon}-\mathit{U}\rVert.$
Similarly $\lVert\mathit{\Delta}-\mathit{D}\rVert=\lVert\mathit{\Upsilon}_{uu}-\mathit{U}_{uu}-\mathit{\Upsilon}_{uy}\mathit{\Upsilon}_{yy}^{-1}\mathit{\Upsilon}_{yu}+\mathit{U}_{uy}\mathit{U}_{yy}^{-1}\mathit{U}_{yu}\rVert\leqslant 4R^{8}\lVert\mathit{\Upsilon}-\mathit{U}\rVert.$ By Schur decomposition, it holds that
$\begin{pmatrix}\mathit{U}_{uu}&\mathit{\mathit{missing}}U_{uy}\\ \mathit{U}_{uy}^{\mathsf{T}}&\mathit{U}_{yy}\end{pmatrix}=\begin{pmatrix}\mathit{I}_{d}&\mathit{U}_{uy}\mathit{U}_{yy}^{-1}\\ 0&\mathit{I}_{K}\end{pmatrix}\begin{pmatrix}\mathit{U}_{uu}-\mathit{U}_{uy}\mathit{U}_{yy}^{-1}\mathit{U}_{uy}^{\mathsf{T}}&0\\ 0&\mathit{U}_{yy}\end{pmatrix}\begin{pmatrix}I_{d}&0\\ \mathit{U}_{yy}^{-1}\mathit{U}_{uy}^{\mathsf{T}}&I_{K}\end{pmatrix}.$ (B.58)
Since $\mathit{D}=\mathit{U}_{uu}-\mathit{U}_{uy}\mathit{U}_{yy}^{-1}\mathit{U}_{uy}^{\mathsf{T}}$, we have by the Courant–Fischer theorem
$\lambda_{\min}(\mathit{U})$ $=\min\left\{\frac{x^{\mathsf{T}}\mathit{U}x}{x^{\mathsf{T}}x}:x\in^{d+K}\setminus\{0\}\right\}\leqslant\min\left\{\frac{x^{\mathsf{T}}\mathit{U}x}{x^{\mathsf{T}}x}:x=\begin{pmatrix}u\\ -\mathit{U}_{yy}^{-1}\mathit{U}_{uy}^{\mathsf{T}}u\end{pmatrix}\text{ and }u\in^{d}\setminus\{0\}\right\}$
$=\min\left\{\frac{u^{\mathsf{T}}Du}{x^{\mathsf{T}}x}:x=\begin{pmatrix}u\\ -\mathit{U}_{yy}^{-1}\mathit{U}_{uy}^{\mathsf{T}}u\end{pmatrix}\text{ and }u\in^{d}\setminus\{0\}\right\}$
$\leqslant\min\left\{\frac{u^{\mathsf{T}}Du}{u^{\mathsf{T}}u}:u\in^{d}\setminus\{0\}\right\}=\lambda_{\min}(D).$
Therefore $\lVert D^{-1}\rVert\leqslant\lVert\mathit{U}^{-1}\rVert\leqslant R^{2}$. Using Lemma 25, we then have
$d_{g}(\mathsf{B}_{j}\pi,\mathsf{B}_{j}p)$ $\leqslant\sqrt{\mathsf{B}_{j}\pi[g^{2}]+\mathsf{B}_{j}p[g^{2}]}\left(3\lVert\mathit{D}^{-1}\mathit{\Delta}-I_{d}\rVert_{\rm F}+\lvert\lambda-\ell\rvert_{\mathit{D}}\right)$
$\leqslant\sqrt{\mathsf{B}_{j}\pi[g^{2}]+\mathsf{B}_{j}p[g^{2}]}\left(3\sqrt{d}\lVert\mathit{\Delta}-\mathit{D}\rVert+\lvert\lambda-\ell\rvert\right)\left(1+\lVert\mathit{D}^{-1}\rVert\right)$
$\leqslant C(R,\kappa_{y})\bigl{(}\lVert\mathit{\Upsilon}-\mathit{U}\rVert+\lvert{\tau}-{t}\rvert\bigr{)},$
for an appropriate constant $C$ depending on $R$ and $\kappa_{y}$. By Lemma 18 it holds that
$\lVert\mathit{\Upsilon}-\mathit{U}\rVert+\bigl{\lvert}{\tau}-{t}\bigr{\rvert}\leqslant\left(\frac{3}{2}+\frac{1}{2}\lvert{\tau}+{t}\rvert\right)\,d_{g}(\pi,p),$
enabling us to conclude. ∎
Appendix C Technical Results for Corollaries 6 and 9
Lemma 30.
Suppose that $\Psi\colon^{d}\to^{d}$ and $h\colon^{d}\to^{K}$ are functions such that Assumption H is satisfied, and let $(\Psi_{n})_{n\in\mathbf{N}}$ and $(h_{n})_{n\in\mathbf{N}}$ be sequences of operators such that
[…]
Let $\mathsf{P}_{n}$ and $\mathsf{Q}_{n}$ denote the maps (1.5) and (1.7) associated with the functions $\Psi_{n}$ and $h_{n}$ and assume that they too satisfy Assumption H. Then
[…]
for all $R\geqslant 1$ for the second statement.
Proof.
For all $\mu\in\mathcal{P}(^{d})$, the probability measures $\mathsf{P}\mu$ and $\mathsf{P}_{n}\mu$ have Lebesgue densities. By the definition in equation (1.5) of $\mathsf{P}$ and Remark 2, it holds that
$d_{g}(\mathsf{P}\mu,\mathsf{P}_{n}\mu)=\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}g(u)\left\lvert\int_{{}^{d}}\Delta_{n}(u,v)\,\mu(\mathrm{d}v)\right\rvert\,\mathrm{d}u,$
where
$\Delta_{n}(u,v):=\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)-\exp\left(-\frac{1}{2}\left\lvert u-\Psi_{n}(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)$
Multiplying and dividing by the density of $\mathcal{N}(0,2\mathit{\Sigma})$ in the outermost integral, we obtain that
$d_{g}(\mathsf{P}\mu,\mathsf{P}_{n}\mu)$ $\leqslant\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}\int_{{}^{d}}\exp\left(\frac{1}{2}\left\lvert u\right\rvert_{2\mathit{\Sigma}}^{2}\right)g(u)\left|\Delta_{n}(u,v)\right|\,\mu(\mathrm{d}v)\,\exp\left(-\frac{1}{2}\left\lvert u\right\rvert_{2\mathit{\Sigma}}^{2}\right)\,\mathrm{d}u$
$\leqslant\sqrt{2^{d}}\sup_{u,v\in^{d}}\left\{\exp\left(\frac{1}{2}\left\lvert u\right\rvert_{2\mathit{\Sigma}}^{2}\right)g(u)\bigl{\lvert}\Delta_{n}(u,v)\bigr{\rvert}\right\}.$ (C.2)
In order to prove the first limit in (C.1), it remains to show that the supremum in (C.2) tends to 0 in the limit as $n\to\infty$. To this end, suppose for contradiction that there was a sequence $(u_{n},v_{n})\in^{d}\times^{d}$ and $\alpha>0$ such that
$\forall n\in\mathbf{N},\qquad\exp\left(\frac{1}{2}\left\lvert u_{n}\right\rvert_{2\mathit{\Sigma}}^{2}\right)g(u_{n})\bigl{\lvert}\Delta_{n}(u_{n},v_{n})\bigr{\rvert}\geqslant\alpha.$ (C.3)
By the assumption that $\Psi_{n}\to\Psi$ in $L^{\infty}(^{d})$, it holds that $\Delta_{n}(u_{n},v_{n})\to 0$, and so it must necessarily hold that $\lvert u_{n}\rvert\to\infty$ for (C.3) to hold. But this leads to a contradiction because, by Young’s inequality, it holds for all $\delta>0$ that
$\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\geqslant\frac{1}{1+\delta}\left\lvert u\right\rvert_{\mathit{\Sigma}}^{2}-\frac{1}{\delta}\left\lvert\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}.$
Therefore, given that $\Psi_{n}\to\Psi$ in $L^{\infty}(^{d})$, the term $\Delta_{n}(u_{n},v_{n})$ in (C.3) decreases to 0 in the limit $\lvert u_{n}\rvert\to\infty$ faster than the other terms grow.
For the second limit in (C.1), fix $\mu\in\mathcal{P}_{R}(^{d})$ and note that, by definition (1.7) of $\mathsf{Q}$,
$d_{g}(\mathsf{Q}\mu,\mathsf{Q}_{n}\mu)$ $\leqslant\frac{1}{\sqrt{(2\pi)^{K}\det\mathit{\Gamma}}}\int_{{}^{d}\times^{K}}\bigl{(}1+\lvert u\rvert^{2}+\lvert y\rvert^{2}\bigr{)}\left|\widetilde{\Delta}_{n}(u,y)\right|\,\mu(\mathrm{d}u)\,\mathrm{d}y$
$\leqslant\frac{1}{\sqrt{(2\pi)^{K}\det\mathit{\Gamma}}}\int_{{}^{d}}\bigl{(}1+\lvert u\rvert^{2}\bigr{)}\left(\int_{{}^{K}}\bigl{(}1+\lvert y\rvert^{2}\bigr{)}\bigl{\lvert}\widetilde{\Delta}_{n}(u,y)\bigr{\rvert}\,\mathrm{d}y\right)\,\mu(\mathrm{d}u),$ (C.4)
where we introduced
$\widetilde{\Delta}_{n}(u,y):=\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y-h_{n}(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)$
The innermost integral in (C.4), which is independent of $\mu$, converges to 0 uniformly in $u\in^{d}$, and so the conclusion follows. ∎
Proposition 31.
Suppose that $\Psi\colon^{d}\to^{d}$ and $h\colon^{d}\to^{K}$ are functions taking constant values, and let $(\Psi_{n})_{n\in\mathbf{N}}$ and $(h_{n})_{n\in\mathbf{N}}$ be sequences of functions such that
[…]
Let $\mathsf{P}_{n}$ and $\mathsf{Q}_{n}$ denote the maps (1.5) and (1.7) (associated with the functions $\Psi_{n}$ and $h_{n}$) and assume that they too satisfy Assumption H. Denote by $(\mu^{n}_{j})_{j\in}$ the true filtering distribution associated with functions $\Psi_{n},h_{n}$. Then, with the same notation as in Lemma 30, it holds that
$\lim_{n\to\infty}\max_{j\in}d_{g}(\mathsf{G}\mathsf{Q}_{n}\mathsf{P}_{n}\mu^{n}_{j},\mathsf{Q}_{n}\mathsf{P}_{n}\mu^{n}_{j})=0.$
Proof.
The conclusion follows from the stronger statement that
$\lim_{n\to\infty}\sup_{\mu\in\mathcal{P}(^{d})}d_{g}(\mathsf{G}\mathsf{Q}_{n}\mathsf{P}_{n}\mu,\mathsf{Q}_{n}\mathsf{P}_{n}\mu)=0.$
Indeed, fix $\mu\in\mathcal{P}(^{d})$. By the triangle inequality, we have that
$d_{g}(\mathsf{G}\mathsf{Q}_{n}\mathsf{P}_{n}\mu,\mathsf{Q}_{n}\mathsf{P}_{n}\mu)\leqslant d_{g}(\mathsf{G}\mathsf{Q}_{n}\mathsf{P}_{n}\mu,\mathsf{G}\mathsf{Q}\mathsf{P}\mu)+d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu)+d_{g}(\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}_{n}\mathsf{P}_{n}\mu).$ (C.5)
Since $\Psi$ and $h$ are constant, the probability measure $\mathsf{Q}\mathsf{P}\mu$ is a Gaussian distribution independent of $\mu$, and so the second term on the right-hand side is zero. For the third term, we have using Lemma 22, that
$d_{g}(\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}_{n}\mathsf{P}_{n}\mu)$ $\leqslant d_{g}(\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}_{n}\mathsf{P}\mu)+d_{g}(\mathsf{Q}_{n}\mathsf{P}\mu,\mathsf{Q}_{n}\mathsf{P}_{n}\mu)$
$\leqslant d_{g}(\mathsf{Q}\nu,\mathsf{Q}_{n}\nu)+\Bigl{(}1+\kappa_{h_{n}}^{2}+\operatorname*{tr}(\mathit{\Gamma})\Bigr{)}d_{g}(\mathsf{P}\mu,\mathsf{P}_{n}\mu),\qquad\nu:=\mathsf{P}\mu.$
Noting that $\nu:=\mathsf{P}\mu$ is a fixed Gaussian measure independent of $\mu$ and using Lemma 30, we deduce that both terms on the right-hand side tend to 0 in the limit $n\to\infty$, uniformly in $\mu\in\mathcal{P}(^{d})$. It remains to show that the first term in (C.5) also tends to 0 as $n\to\infty$ uniformly in $\mu\in\mathcal{P}(^{d})$. In view of the moment bounds in Lemma 16, there exist $R\geqslant 1$ and $N\in\mathbf{N}$ such that
$\forall n\geqslant N,\qquad\forall\mu\in\mathcal{P}(^{d}),\qquad\mathsf{Q}_{n}\mathsf{P}_{n}\mu\in\mathcal{P}_{R}(^{d})\qquad\text{and}\qquad\mathsf{Q}\mathsf{P}\mu\in\mathcal{P}_{R}(^{d}).$
Therefore, by the local Lipschitz continuity of $\mathsf{G}$ established in Lemma 26, and the uniform-in-$\mu$ convergence to 0 of the third term in (C.5) that we already proved, it holds that
$\sup_{\mu\in\mathcal{P}(^{d})}d_{g}(\mathsf{G}\mathsf{Q}_{n}\mathsf{P}_{n}\mu,\mathsf{G}\mathsf{Q}\mathsf{P}\mu)\xrightarrow[n\to\infty]{}0,$
which concludes the proof. ∎